### v3

# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmarkthanks: This paper has been accepted by CVPR 2025

Abstract
AI-generated faces have enriched human life, such as entertainment, education, and art. However, they also pose misuse risks. Therefore, detecting AI-generated faces becomes crucial, yet current detectors show biased performance across different demographic groups. Mitigating biases can be done by designing algorithmic fairness methods, which usually require demographically annotated face datasets for model training. However, no existing dataset encompasses both demographic attributes and diverse generative methods simultaneously, which hinders the development of fair detectors for AI-generated faces. In this work, we introduce the AI-Face dataset, the first million-scale demographically annotated AI-generated face image dataset, including real faces, faces from deepfake videos, and faces generated by Generative Adversarial Networks and Diffusion Models. Based on this dataset, we conduct the first comprehensive fairness benchmark to assess various AI face detectors and provide valuable insights and findings to promote the future fair design of AI face detectors. Our AI-Face dataset and benchmark code are publicly available at https://github.com/Purdue-M2/AI-Face-FairnessBench.
footnotetext: Corresponding author
1 Introduction
AI-generated faces are created using sophisticated AI technologies that are visually difficult to discern from real ones [1]. They can be summarized into three categories: deepfake videos [2] created by typically using Variational Autoencoders (VAEs) [3, 4], faces generated from Generative Adversarial Networks (GANs) [5, 6, 7, 8], and Diffusion Models (DMs) [9]. These technologies have significantly advanced the realism and controllability of synthetic facial representations. Generated faces can enrich media and increase creativity [10]. However, they also carry significant risks of misuse. For example, during the 2024 United States presidential election, fake face images of Donald Trump surrounded by groups of black people smiling and laughing to encourage African Americans to vote Republican are spreading online [11]. This could distort public opinion and erode people’s trust in media [12, 13], necessitating the detection of AI-generated faces for their ethical use.
Figure 1: Comparison between AI-Face and other datasets in terms of demographic annotation, generation category, and the number of generation methods. ‘DF’, ‘GAN’, and ‘DM’ stand for Deepfake Videos, Generative Adversarial Networks, and Diffusion Models.
  Face Images Generation Category Demographic Annotation
Dataset Year #Real #Fake Deepfake Videos GAN DM #Generation Methods Source of Real Images Skin Tone Gender Age
DF-1.0 [14] 2020 2.9M 14.7M 1 Self-Recording
DeePhy [15] 2022 1K 50.4K 3 YouTube
DF-Platter [16] 2023 392.3K 653.4K 3 YouTube
GenData [17] 2023 - 20K 3 CelebA [18]
A-FF++ [19] 2024 29.8K 149.1K 5 YouTube
A-DFD [19] 2024 10.8K 89.6K 5 Self-Recording
A-DFDC [19] 2024 54.5K 52.6K 8 Self-Recording
A-Celeb-DF-v2 [19] 2024 26.3K 166.5K 1 Self-Recording
A-DF-1.0 [19] 2024 870.3K 321.5K 1 Self-Recording
AI-Face 2025 400K 1.2M 37
FFHQ [6], IMDB-WIKI [20],
real from FF++ [2], DFDC [21],
DFD [22],Celeb-DF-v2 [23]
Table 1: Quantitative comparison of existing datasets with ours on demographically annotated AI-generated faces.
However, one major issue existing in current AI face detectors [24, 25, 26, 27] is biased detection (i.e., unfair detection performance among demographic groups [28, 19, 29, 30]). Mitigating biases can be done by designing algorithmic fairness methods, but they usually require demographically annotated face datasets for model training. For example, works like  [29, 30] have made efforts to enhance fairness in the detection based on A-FF++ [19] and A-DFD [19]. However, both datasets are limited to containing only faces from deepfake videos, which could cause the trained models not to be applicable for fairly detecting faces generated by GANs and DMs. While some datasets (e.g., GenData [17], DF40 [31]) include GAN and DM faces, they either lack demographic annotations or provide only limited demographic attributes. Most importantly, no existing dataset offers sufficient diversity in generation methods while also providing demographic labels. A comparison of existing datasets is shown in Fig. 1. These limitations of existing datasets hamper the development of fair technologies for detecting AI-generated faces.
Moreover, benchmarking fairness provides a direct method to uncover prevalent and unique fairness issues in recent AI-generated face detection. However, there is a lack of a comprehensive benchmark to estimate the fairness of existing AI face detectors. Existing benchmarks [32, 33, 34, 35] primarily assess utility, neglecting systematic fairness evaluation. Two studies [28, 36] do evaluate fairness in detection models, but their examination is based on a few outdated detectors. Furthermore, detectors’ fairness reliability (e.g., robustness with test set post-processing, fairness generalization) has not been assessed. The absence of a comprehensive fairness benchmark impedes a thorough understanding of the fairness behaviors of recent AI face detectors and obscures the research path for detector fairness guarantees.
In this work, we build the first million-scale demographically annotated AI-generated face image dataset: AI-Face. The face images are collected from various public datasets, including the real faces that are usually used to train AI face generators, faces from deepfake videos, and faces generated by GANs and DMs. Each face is demographically annotated by our designed measurement method and Contrastive Language-Image Pretraining (CLIP) [37]-based lightweight annotator. Next, we conduct the first comprehensive fairness benchmark on our dataset to estimate the fairness performance of 12 representative detectors coming from four model types. Our benchmark exposes common and unique fairness challenges in recent AI face detectors, providing essential insights that can guide and enhance the future design of fair AI face detectors. Our contributions are as follows:
  • We build the first million-scale demographically annotated AI-generated face dataset by leveraging our designed measure and developed lightweight annotator.
  • We conduct the first comprehensive fairness benchmark of AI-generated face detectors, providing an extensive fairness assessment of current representative detectors.
  • Based on our experiments, we summarize the unsolved questions and offer valuable insights within this research field, setting the stage for future investigations.
2 Background and Motivation
AI-generated Faces and Biased Detection. AI-generated face images, created by advanced AI technologies, are visually difficult to discern from real ones. They can be summarized into three categories: 1) Deepfake Videos. Initiated in 2017 [13], these use face-swapping and face-reenactment techniques with a variational autoencoder to replace a face in a target video with one from a source [3, 4]. Note that our paper focuses solely on images extracted from videos. 2) GAN-generated Faces. Post-2017, Generative Adversarial Networks (GANs) [38] like StyleGANs [6, 7, 8] have significantly improved generated face realism. 3) DM-generated Faces. Diffusion models (DMs), emerging in 2021, generate detailed faces from textual descriptions and offer greater controllability. Tools like Midjourney [39] and DALLE2 [40] facilitate customized face generation. While these AI-generated faces can enhance visual media and creativity [10], they also pose risks, such as being misused in social media profiles [41, 42]. Therefore, numerous studies focus on detecting AI-generated faces [24, 25, 26, 27], but current detectors often show performance disparities among demographic groups [28, 19, 29, 30]. This bias can lead to unfair targeting or exclusion, undermining trust in detection models. Recent efforts [29, 30] aim to enhance fairness in deepfake detection but mainly address deepfake videos, overlooking biases in detecting GAN- and DM-generated faces.
The Existing Datasets. Current AI-generated facial datasets with demographic annotations are limited in size, generation categories, methods, and annotations, as illustrated in Table 1. For instance, A-FF++, A-DFD, A-DFDC, and A-Celeb-DF-v2 [19] are deepfake video datasets with fewer than one million images. Datasets like DF-1.0 [14] and DF-Platter [16] lack various demographic annotations. Additionally, existing datasets offer limited generation methods. These limitations hinder the development of fair AI face detectors, motivating us to build a million-scale demographically annotated AI-Face dataset.
Benchmark for Detecting AI-generated Faces. Benchmarks are essential for evaluating AI-generated face detectors under standardized conditions. Existing benchmarks, as shown in Table 2, mainly focus on detectors’ utility, often overlooking fairness [32, 33, 34, 35, 31]. Loc et al. [28] and CCv1 [36] examined detector fairness. However, their study did not have an analysis on DM-generated faces and only measured bias between groups in basic scenarios without considering fairness reliability under real-world variations and transformations. This motivates us to conduct a comprehensive benchmark to evaluate AI face detectors’ fairness.
The Definition of Demographic Categories. Demography-related labels are highly salient to measuring bias. Following prior works [43, 44, 45, 36, 46, 47], we will focus on three key demographic categories: Skin Tone, Gender, and Age, in this work. For skin tone, this vital attribute spans a range from pale to dark. We use the Monk Skin Tone scale [48], specifically designed for computer vision applications. For gender, we adopt binary categories (i.e., Male and Female), following practices by many governments [49, 50] and facial recognition research [45, 51, 52], based on sex at birth. For age, using definitions from the United Nations [53] and Statistics Canada [54], we define five age groups: Child (0-14), Youth (15-24), Adult (25-44), Middle-age Adult (45-64), and Senior (65+). More discussion is in Appendix A.
3 AI-Face Dataset
This section outlines the process of building our demographically annotated AI-Face dataset (see Fig. 2), along with its statistics and annotation quality assessment.
  Category Scope of Benchmark
Fairness
Existing Benchmarks Year Deepfake Videos GAN DM Utility General Reliability
Loc et al. [28] 2021
CCv1 [36] 2021
DeepfakeBench [34] 2023
CDDB [32] 2023
Lin et al. [33] 2024
Le et al. [35] 2024
DF40 [31] 2024
Ours 2025
Table 2: Comparison of existing AI-generated face detection benchmarks and ours. Fairness ‘General’ means fairness evaluation under default/basic settings. Fairness ‘Reliability’ measures fairness consistency across dynamic scenarios (e.g., post-processing).
Figure 2: Generation pipeline of our Demographically Annotated AI-Face Dataset. First, we collect and filter face images from Deepfake Videos, GAN-generated faces, and DM-generated faces found in public datasets. Second, we perform skin tone, gender, and age annotation generation. Skin tone is estimated by combining facial landmark detection with color analysis to generate the corresponding annotation. For gender and age, we develop annotators trained on the IMDB-WIKI dataset [20], then use them to predict attributes for each image.
3.1 Data Collection
We build our AI-Face dataset by collecting and integrating public real and AI-generated face images sourced from academic publications, GitHub repositories, and commercial tools. We strictly adhere to the license agreements of all datasets to ensure that they allow inclusion in our datasets and secondary use for training and testing. More details are in Appendix B.1. Specifically, the fake face images in our dataset originate from 4 Deepfake Video datasets (i.e., FF++ [2], DFDC [21], DFD [22], and Celeb-DF-v2 [23]), generated by 10 GAN models (i.e., AttGAN [55], MMDGAN [56], StarGAN [55], StyleGANs [55, 57, 58], MSGGAN [56], ProGAN [59], STGAN [56], and VQGAN [60]), and 8 DM models (i.e., DALLE2 [61], IF [61], Midjourney [61], DCFace [62], Latent Diffusion [63], Palette [64], Stable Diffusion v1.5 [65], Stable Diffusion Inpainting [65]). This constitutes a total of 1,245,660 fake face images in our dataset. We include 6 real source datasets (i.e., FFHQ [6], IMDB-WIKI [20], and real images from FF++ [2], DFDC [21], DFD [22], and Celeb-DF-v2 [23]). All of them are usually used as a training set for generative models to generate fake face images. This constitutes a total of 400,885 real face images in our dataset. In general, our dataset contains 28 subsets and 37 generation methods (i.e., 5 in FF++, 5 in DFD, 8 in DFDC, 1 in Celeb-DF-v2, 10 GANs, and 8 DMs). For all images, we use RetinaFace [66] for detecting and cropping faces.
3.2 Annotation Generation
3.2.1 Skin Tone Annotation Generation
Skin tone is typically measured using an intuitive approach [67, 68], without requiring a predictive model. Inspired by [67], we developed a method to estimate skin tone using the Monk Skin Tone (MST) Scale [48] (including 10-shade scales: Tone 1 to 10) by combining facial landmark detection with color analysis. Specifically, utilizing Mediapipe’s FaceMesh [69] for precise facial landmark localization, we isolate skin regions while excluding non-skin areas such as the eyes and mouth. Based on the detected landmarks, we generate a mask to extract skin pixels from the facial area. These pixels are then subjected to K-Means clustering [70] (we use K$=3$ in practice) to identify the dominant skin color within the region of interest. The top-1 largest color cluster is mapped to the closest tone in the MST Scale by calculating the Euclidean distance between the cluster centroid and the MST reference colors in RGB space.
3.2.2 Gender and Age Annotation Generation
For generating gender and age annotations, the existing online software (e.g., Face++ [71]) and open-source tools (e.g., InsightFace [72]) can be used for the prediction. However, they fall short in our task due to two reasons: 1) They are mostly designed for face recognition and trained on datasets of real face images but lack generalization capability for annotating AI-generated face images. 2) Their use may introduce bias into our dataset, as they are typically designed and trained without careful consideration of bias and imbalance in the training set. See Appendix  B.3 for our experimental study on these tools. To this end, we have to develop our specific annotators to predict gender and age annotations for each image in our dataset.
[…]
Annotator Architecture. We build a lightweight annotator based on the CLIP [37] foundation model by leveraging its strong zero-shot and few-shot learning capabilities. Specifically, our annotator employs a frozen pre-trained CLIP ViT L/14 [73] as a feature extractor $\mathbf{E}$ followed by a trainable classifier parameterized by $\theta$, which contains 3-layer Multilayer Perceptron (MLP) $\mathbf{M}$ and a classification head $h$.
Learning Objective. Aware that neural networks can perform poorly when the training dataset suffers from class-imbalance [74] and CLIP is not free from demographic bias [75, 76, 77], we introduce an imbalance loss and fairness loss to address these challenges in the annotator training. Specifically, for image $X_{i}$, its feature $f_{i}$ is obtained through $f_{i}\!\!=\!\!\mathbf{M}(\mathbf{E}(X_{i}))$. Next, two losses are detailed below.
Imbalance Loss: To mitigate the impact of imbalance data, we use Vector Scaling [78] loss, which is a re-weighting method for training models on the imbalanced data with distribution shifts and can be expressed as
$L_{imb}=\frac{1}{n}\sum_{i=1}^{n}-u_{A_{i}}\log\frac{e^{\zeta_{A_{i}}h(f_{i})_{A_{i}}+\Delta_{A_{i}}}}{\sum_{A\in\mathcal{A}}e^{\zeta_{A}h(f_{i})_{A}+\Delta_{A}}},\vspace{-1mm}$
where $u_{A_{i}}$ is the weighting factor for attribute $A_{i}$. $h(f_{i})_{A_{i}}$ is the predict logit on $A_{i}$. $\zeta_{A_{i}}$ is the multiplicative logit scaling factor, calculated as the inverse of $A_{i}$’s frequency. $\Delta_{A_{i}}$ is the additive logit scaling factor, calculated as the log of $A_{i}$ probabilities. More details about them are in appendix B.4.
Fairness Loss: We introduce a fairness loss to minimize the disparity between the distribution $\mathcal{D}^{f}$ of $f$ and the conditional distribution $\mathcal{D}^{f_{A}}$ of $f$ on attribute $A\in\mathcal{A}$. Specifically, we follow [79, 80] to minimize the summation of the following Sinkhorn distance between these two distributions:
$L_{fair}=\sum_{A\in\mathcal{A}}\inf_{\gamma\in\Gamma(\mathcal{D}^{f},\mathcal{D}^{f_{A}})}\big{\{}\mathbb{E}_{X\sim\gamma}[c(p,q)]+\alpha H(\gamma|\mu\otimes\nu)\big{\}},$
where $\Gamma(\mathcal{D}^{f},\mathcal{D}^{f_{A}})$ is the set of joint distributions based on $\mathcal{D}^{f}$ and $\mathcal{D}^{f_{A}}$. Let $p$ and $q$ be the points from $\mathcal{D}^{f}$ and $\mathcal{D}^{f_{A}}$, respectively. Then, $c(p,q)$ represents the transport cost [80]. Let $\mu$ and $\nu$ be the reference measures from the set of measures on $f$. Then, $H(\gamma|\mu\otimes\nu)$ represents the relative entropy of $\gamma$ with respect to the product measure $\mu\otimes\nu$. $\alpha\geq 0$ is a regularization hyperparameter. In practice, we use the empirical form of $L_{fair}$.
Total Loss: Therefore, the final learning objective becomes $\mathcal{L}(\theta)=L_{imb}+\lambda L_{fair}$, where $\lambda$ is a hyperparameter.
Figure 3: Distribution of face images of the AI-Face dataset. The figure shows the (a) subset distribution and the demographic distribution for (b) skin tone, (c) gender, and (d) gender. The outer rings in (b), (c), and (d) represent the proportion of groups within each attribute category, while the inner rings indicate the distribution of fake (F) and real (R) images within those groups.
Train. Traditional optimization methods like stochastic gradient descent can lead to poor model generalization due to sharp loss landscapes with multiple local and global minima. To address this, we use Sharpness-Aware Minimization (SAM) [81] to enhance our annotator’s generalization by flattening the loss landscape. Specifically, flattening is attained by determining the optimal $\epsilon^{*}$ for perturbing model parameters $\theta$ to maximize the loss, formulated as: $\epsilon^{*}=\arg\max_{\|\epsilon\|_{2}\leq\beta}{\mathcal{L}}\textbf{(}\theta+\epsilon\textbf{)}\approx\arg\max_{\|\epsilon\|_{2}\leq\beta}\epsilon^{\top}\nabla_{\theta}\mathcal{L}=\beta\texttt{sign}(\nabla_{\theta}\mathcal{L})$, where $\beta$ controls the perturbation magnitude. The approximation is based on the first-order Taylor expansion with assuming $\epsilon$ is small. The final equation is obtained by solving a dual norm problem, where sign represents a sign function and $\nabla_{\theta}\mathcal{L}$ being the gradient of $\mathcal{L}$ with respect to $\theta$. As a result, the model parameters are updated by solving: $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}$.
Inference. We use the trained annotators to predict demographic labels for each image in AI-Face dataset, except for those from IMDB-WIKI, which already contain true labels.
3.3 Dataset Statistics
Fig. 3 illustrates the subset distribution and demographic attributes of the AI-Face dataset. The dataset contains approximately three times more generated images than real images, with diffusion model-generated images constituting the majority. In terms of demographic attributes, the majorities in skin tone are Tone 5 (31.14%) and Tone 6 (35.16%). The lightest skin tones (Tones 1-3) are underrepresented, comprising only 0.97% of the dataset. The dataset is relatively balanced across gender. Adult (25-44) (49.67%) is the predominant representation in age groups.
3.4 Annotation Quality Assessment
To assess the quality of demographic annotations in our AI-Face dataset, we conducted a user study. Three participants label the demographic attributes for the given images (the details of labeling activities are in appendix B.5), with the final ground truth determined by majority vote. We then compare our annotations with those in A-FF++, A-DFDC, A-CelebDF-V2, and A-DFD datasets. Specifically, we perform two assessments: 1) Strategic comparison: We select 1,000 images from A-FF++ and A-DFDC that have different annotations from AI-Face. These images likely represent challenging cases. 2) Random comparison: We randomly sampled 1,000 images from A-Celeb-DF-V2 and A-DFD. Due to the limited age classes in these datasets, only gender was evaluated. The results, presented in Table 3, demonstrate the high correctness of the AI-Face annotations and their superior quality compared to the annotations of other datasets. For example, our annotation quality (ACC) surpasses those in A-FF++ by 78.714% on gender and 48.000% on age.
4 Fairness Benchmark Settings
This section demonstrates the fairness benchmark settings for detection methods and evaluation metrics on AI-Face (80%/20%: Train/Test). More settings are in Appendix C.1.
Detection Methods. Our benchmark has implemented 12 detectors. The methodologies cover a spectrum that is specifically tailored to detect AI-generated faces from Deepfake Videos, GANs, and DMs. They can be classified into four types: Naive detectors: refer to backbone models that can be directly utilized as the detector for binary classification, including CNN-based (i.e., Xception [82], EfficientB4 [83]) and transformer-based (i.e., ViT-B/16 [84]). Frequency-based: explore the frequency domain for forgery detection (i.e., F3Net [85], SPSL [86], SRM [87]). Spatial-based: focus on mining spatial characteristics (e.g., texture) within images for detection (i.e., UCF [26], UnivFD [88], CORE [89]). Fairness-enhanced: focus on improving fairness in AI-generated face detection by designing specific algorithms (i.e., DAW-FDD [29], DAG-FDD [29], PG-FDD [30]).
  Gender Age
Evaluation Type Dataset ACC Precision Recall ACC Precision Recall
A-FF++ 8.143 17.583 5.966 37.700 39.459 45.381
AI-Face 86.857 74.404 77.367 85.700 74.024 63.751
A-DFDC 21.600 28.604 23.082 33.400 38.011 40.165
Strategic AI-Face 91.700 92.129 83.448 77.000 76.184 62.646
A-Celeb-DF-V2 89.628 90.626 90.494 -
AI-Face 91.206 91.474 91.767 -
A-DFD 70.900 71.686 74.435 -
Random AI-Face 92.300 91.060 91.727 -
Table 3: Annotation quality assessment results (%) for A-FF++, A-DFDC, A-Celeb-DF-V2, A-DFD, and our AI-Face. ACC: Accuracy.
  Model Type
Naive Frequency Spatial Fairness-enhanced
Measure Attribute Metric
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
$F_{MEO}$ 8.836 8.300 6.264 19.938 8.055 10.002 17.325 2.577 10.779 14.118 6.551 6.465
$F_{DP}$ 9.751 6.184 7.728 12.876 9.379 10.897 12.581 8.556 10.317 10.706 8.617 9.746
$F_{OAE}$ 1.271 4.377 2.168 2.818 1.135 0.915 1.883 2.748 1.332 1.667 1.388 0.882
Skin Tone $F_{EO}$ 12.132 11.062 8.813 23.708 9.789 14.239 21.92 5.536 13.069 16.604 7.383 9.115
$F_{MEO}$ 3.975 5.385 5.104 4.717 4.411 6.271 5.074 4.503 5.795 5.510 5.910 3.190
$F_{DP}$ 1.691 1.725 1.344 1.864 1.827 1.957 1.736 1.190 2.154 2.015 2.151 1.252
$F_{OAE}$ 0.975 1.487 1.803 1.129 1.037 1.772 1.451 1.622 1.389 1.325 1.420 1.071
Gender $F_{EO}$ 4.143 5.863 6.031 4.870 4.534 6.78 5.510 5.408 5.931 5.696 6.066 3.702
$F_{MEO}$ 27.883 6.796 14.937 38.801 27.614 24.843 47.500 5.436 33.882 45.466 15.229 14.804
$F_{DP}$ 10.905 11.849 11.839 14.906 11.232 11.570 17.049 15.249 12.564 14.106 9.633 10.467
$F_{OAE}$ 7.265 2.856 6.838 10.116 7.270 6.524 11.652 3.793 8.760 11.878 5.533 5.009
Age $F_{EO}$ 42.216 10.300 30.795 55.032 40.943 38.528 67.545 14.148 48.729 64.384 30.182 29.585
$F_{MEO}$ 10.505 17.586 9.384 21.369 10.379 15.142 20.134 6.119 15.34 16.565 12.178 9.578
$F_{DP}$ 14.511 8.607 11.535 17.175 13.259 15.186 17.03 14.026 14.301 14.088 11.705 14.697
$F_{OAE}$ 2.536 8.461 4.928 4.870 2.464 3.998 3.536 6.287 2.775 3.547 4.035 3.062
Intersection $F_{EO}$ 24.315 25.114 27.443 47.783 21.679 30.112 43.376 20.255 28.84 33.122 26.295 18.348
Fairness(%)↓ Individual $F_{IND}$ 10.338 25.742 0.022 1.872 2.518 7.621 0.767 3.523 0.041 3.772 0.901 0.780
AUC↑ 98.583 98.611 98.69 98.714 98.747 97.936 98.082 98.192 98.579 97.811 98.771 99.172
ACC↑ 96.308 94.203 94.472 95.719 96.346 95.092 95.151 93.651 96.224 95.426 95.722 96.174
AP↑ 99.350 99.542 99.571 99.453 99.356 99.172 99.273 99.400 99.360 99.015 99.498 99.694
EER↓ 5.149 6.689 6.372 5.256 4.371 6.483 7.708 7.633 5.145 7.063 5.499 4.961
Utility(%) - FPR↓ 12.961 20.066 16.426 14.679 13.661 15.746 13.646 18.550 13.410 16.670 14.844 10.971
Training Time / Epoch 1h15min 2h25min 2h40min 1h18min 1h20min 3h10min 5h05min 4h 1h16min 1h25min 1h17min 7h20min
Table 4: Overall performance comparison of difference methods on the AI-Face dataset. The best performance is shown in bold.
Evaluation Metrics. To provide a comprehensive benchmarking, we consider 5 fairness metrics commonly used in fairness community [90, 91, 92, 93, 94] and 5 widely used utility metrics. For fairness metrics, we consider Demographic Parity ($F_{DP}$[90, 91], Max Equalized Odds ($F_{MEO}$[93], Equal Odds ($F_{EO}$[92], and Overall Accuracy Equality ($F_{OAE}$[93] for evaluating group (e.g., gender) and intersectional (e.g., individuals of a specific gender and simultaneously a specific skin tone) fairness. In experiments, the intersectional groups are Female-Light (F-L), Female-Medium (F-M), Female-Dark (Dark), Male-Light (M-L), Male-Medium (M-M), and Male-Dark (M-D), where we group 10 categories of skin tones into Light (Tone 1-3), Medium (Tone 4-6), and Dark (Tone 7-10) for simplicity according to [95]. We also use individual fairness ($F_{IND}$[94, 96] (i.e., similar individuals should have similar predicted outcomes) for estimation. For utility metrics, we employ the Area Under the ROC Curve (AUC), Accuracy (ACC), Average Precision (AP), Equal Error Rate (EER), and False Positive Rate (FPR).
5 Results and Analysis
In this section, we estimate the existing AI-generated image detectors’ fairness performance alongside their utility on our AI-Face Dataset. More results can be found in Appendix D.
5.1 General Fairness Comparison
Overall Performance. Table 4 reports the overall performance on our AI-Face test set. Our observations are: 1) Fairness-Enhanced Models (specifically PG-FDD [30]) are the most effective in achieving both high fairness and utility, underscoring the effectiveness of specialized fairness-enhancement techniques in mitigating demographic biases. 2) UnivFD [88], based on the CLIP backbone [73], also achieves commendable fairness, suggesting that foundation models equipped with fairness-focused enhancements could be a promising direction for developing fairer detectors. 3) Naive detectors, such as EfficientB4 [83], trained on large, diverse datasets (e.g., our AI-Face) can achieve competitive fairness and utility, highlighting the potential of fairness improvements by choosing specific architecture. 4) 10 out of 12 detectors have an AUC higher than 98%, demonstrating our AI-Face dataset is significant for training AI-face detectors in resulting high utility. 5) PG-FDD demonstrates superior performance but has a long training time, which can be explored and addressed in the future.
Figure 4: Visualization of the intersectional $F_{EO}$ (%) and AUC (%) of detectors on different subsets. The smaller $F_{EO}$ polygon area represents better fairness. The larger AUC area means better utility.
Figure 5: FPR(%) of each intersectional subgroup The dashline represents the lowest FPR on Female-Light (F-L) subgroup.
Performance on Different Subsets. 1) Fig. 4 demonstrates the intersectional $F_{EO}$ and AUC performance of detectors on each test subset. We observe that the fairness performance varies a lot among different generative methods for every detector. The largest bias on most detectors comes from detecting face images generated by diffusion models. 2) DAG-FDD [29] and SRM [87] demonstrate the most consistent fairness across subsets, indicating a robust handling of bias introduced by different generative methods. 3) Moreover, the stable utility demonstrates our dataset’s expansiveness and diversity, enabling effective training to detect AI-generated faces from various generative techniques.
Performance on Different Subgroups. We conduct an analysis of all detectors on intersectional subgroups. 1) As shown in Fig. 5, facial images with lighter skin tone are more often misclassified as fake, likely due to the underrepresentation of lighter tones (Tone 1-3) in our dataset (see Fig. 3 (b)). This suggests detectors tend to show higher error rates for minority groups. 2) Although gender representation is relatively balanced (see Fig. 3 (c)) in our dataset, the detectors consistently exhibit higher false positive rates for female subgroups, indicating a persistent gender-based bias.
5.2 Fairness Reliability Assessment
Fairness Robustness Evaluation. We apply 6 post-processing methods: Random Crop (RC) [97], Rotation (RT) [34], Brightness Contrast (BC) [34], Hue Saturation Value (HSV) [34], Gaussian Blur (GB) [34], and JEPG Compression (JC) [98] to the test images. Fig. 6 shows each detector’s intersectional $F_{EO}$ and AUC performance changes after using post-processing. Our observations are: 1) These impairments tend to wash out forensic traces, so that detectors have evident performance degradation. 2) Post-processing does not always cause detectors more bias (e.g., UCF, UnivFD, CORE, DAW-FDD have better fairness after rotation), though they hurt the utility. 3) Fairness-enhanced detectors struggle to maintain fairness when images undergo post-processing. 4) Spatial detectors have better fairness robustness compared with other model types.
Figure 6: Performance ratio after vs. before post-processing. Points closer to 1.0 (i.e., no post-processing) indicate better robustness.
  Dataset
CCv2 [99] DF-Platter [16] GenData [17]
Fairness(%)↓ Utility(%)↑ Fairness(%)↓ Utility(%)↑ Fairness(%)↓ Utility(%)↑
Model Type Detector $F_{OAE}$ ACC $F_{OAE}$ $F_{EO}$ AUC $F_{OAE}$ $F_{EO}$ AUC
Xception 1.006(+0.031) 86.465(-9.843) 6.836(+5.861) 9.789(+5.646) 81.273(-17.310) 2.539(+1.564) 13.487(+9.344) 96.971(-1.612)
EfficientB4 4.077(+0.259) 82.980(-11.223) 8.786(+7.299) 12.370(+6.507) 67.694(-30.917) 3.304(+1.817) 1.995(-3.686) 93.213(-5.398)
Naive ViT-B/16 2.167(+0.364) 81.489(-12.983) 0.015(-1.788) 12.373(+6.342) 76.050(-22.640) 3.164(+1.361) 9.610(+3.579) 88.253(-10.437)
F3Net 5.743(+4.614) 87.867(-7.852) 3.521(+2.392) 6.445(+1.575) 85.112(-13.602) 1.188(+0.059) 16.306(+11.436) 91.603(-7.111)
SPSL 0.601(-0.436) 80.006(-16.340) 5.109(+4.072) 7.842(+3.308) 82.175(-16.572) 1.385(+0.348) 9.261(+4.272) 98.838(+0.091)
Frequency SRM 7.000(+5.228) 79.768(-15.324) 3.823(+2.051) 6.567(-0.213) 66.401(-31.535) 3.281(+1.509) 7.907(+1.127) 90.049(-7.887)
UCF 2.169(+0.718) 93.009(-2.142) 8.687(+7.236) 17.068(+11.558) 80.821(-17.261) 3.513(+2.062) 10.529(+5.019) 87.778(-10.304)
UnivFD 7.625(+6.003) 67.983(-25.668) 4.540(+2.918) 9.950(+4.542) 76.443(-21.749) 1.645(+0.023) 3.848(-1.560) 94.418(-3.774)
Spatial CORE 4.410(+3.021) 83.328(-12.896) 7.741(+6.352) 17.348(+11.417) 77.226(-21.353) 3.759(+2.370) 23.289(+17.358) 98.408(-0.171)
DAW-FDD 4.726(+3.401) 84.685(-10.741) 5.536(+4.211) 13.667(+7.791) 81.807(-16.004) 1.443(+0.118) 10.228(+4.532) 97.854(+0.043)
DAG-FDD 2.364(+0.944) 83.918(-11.804) 3.064(+1.644) 22.203(+16.137) 75.206(-23.565) 0.714(-0.706) 10.332(+4.266) 92.108(-6.663)
Fairness- enhanced PG-FDD 1.513(+0.442) 92.852(-3.322) 4.565(+3.494) 9.717(+6.015) 85.271(-13.901) 3.063(+1.992) 9.479(+5.777) 93.329(-5.843)
Table 5: Fairness generalization results based on the gender attribute. The smallest performance changes (in parentheses) and the best performance are in red and in bold, respectively. Only $F_{OAE}$ fairness metric and ACC metric are used in CCv2 due to all samples are real.
Figure 7: Impact of the training set size (Left) and the ratio of real and fake (Right) on detectors’ intersectional $F_{EO}$(%) and AUC (%).
Fairness Generalization Evaluation. To evaluate detectors’ fairness generalization capability, we test them on Casual Conversations v2 (CCv2) [99], DF-Platter [16], and GenData [17], none of which are part of AI-Face. Notably, CCv2 is a dataset that contains only real face images with demographic annotations (e.g., gender) self-reported by the participants. Results on gender attribute in Table 5 show that: 1) Even well-designed detectors that focus on improving utility or fairness generalization (e.g., UCF, PG-FDD) struggle to achieve consistently superior performance across different dataset domains. This highlights the remaining fairness generalization issue. 2) DAW-FDD and PG-PDD are two fairness-enhanced detectors that require accessing demographic information during training, but their fairness does not encounter a drastic drop when evaluating on CCv2. This reflects the high accuracy of the annotations in our AI-face.
Effect of Training Set Size. We randomly sample 20%, 40%, 60%, and 80% of each training subset from AI-Face to assess the impact of training size on performance. Key observations from Fig. 7 (Left): 1) Among all detectors, UnivFD demonstrates the most stable fairness and utility performance as the training dataset size changes, likely due to its fixed CLIP backbone. 2) Increasing the training dataset size generally improves model utility, but this pattern does not extend to fairness metrics. In fact, certain detectors such as F3Net and UCF exhibit worsening fairness as the training size reaches its maximum. This suggests that more training data does not necessarily lead to fairer detectors.
Effect of the Ratio of Real and Fake. To examine how training real-to-fake sample ratios affect detector performance, we set the ratios at 1:10, 1:1, and 10:1 while keeping the total sample count constant. Experimental results in Fig. 7 (Right) show: 1) Most detectors’ fairness improves as real sample representation increases. Probably because increasing real and reducing fake samples helps detectors reduce overfitting to artifacts specific to fake samples. This makes it easier for detectors to distinguish real from fake, even for underrepresented groups, thereby enhancing fairness. 2) Most detectors achieve the highest AUC with balanced data.
5.3 Discussion
According to the above experiments, we summarize the unsolved fairness problems in recent detectors: 1) Detectors’ fairness is unstable when detecting face images generated by different generative methods, indicating a future direction for enhancing fairness stability since new generative models continue to emerge. 2) Even though fairness-enhanced detectors exhibit small overall fairness metrics, they still show biased detection towards minority groups. Future studies should be more cautious when designing fair detectors to ensure balanced performance across all demographic groups. 3) There is currently no reliable detector, as all detectors experience severe large performance degradation under image post-processing and cross-domain evaluation. Future studies should aim to develop a unified framework that ensures fairness, robustness, and generalization, as these three characteristics are essential for creating a reliable detector. Moreover, integrating foundation models (e.g., CLIP) into detector design may help mitigate bias.
6 Conclusion
This work presents the first demographically annotated million-scale AI-Face dataset, serving as a pivotal foundation for addressing the urgent need for developing fair AI face detectors. Based on this dataset, we conduct the first comprehensive fairness benchmark, shedding light on the fairness performance and challenges of current representative AI face detectors. Our findings can inspire and guide researchers in refining current models and exploring new methods to mitigate bias. Limitation and Future Work: One limitation is that our dataset’s annotations are algorithmically generated, so they may lack 100% accuracy. This challenge is difficult to resolve, as demographic attributes for most AI-generated faces are often too ambiguous to predict and do not map to real-world individuals. We plan to enhance annotation quality through human labeling in the future. We also plan to extend our fairness benchmark to evaluate large language models like LLaMA2 [100] and GPT4 [101] for detecting AI faces. Social Impact: Malicious users could misuse AI-generated face images from our dataset to create fake social media profiles and spread misinformation. To mitigate this risk, only users who submit a signed end-user license agreement will be granted access to our dataset.
Ethics Statement
Our dataset collection and annotation generation are approved by Purdue’s Institutional Review Board. The dataset is only for research purposes. All data included in this work are sourced from publicly available datasets, and we strictly comply with each dataset’s license agreement to ensure lawful inclusion and permissible secondary use for training and testing. All collected data and their associated licenses are mentioned in the Datasheet of AI-Face in Appendix E. Our annotation processes prioritize ethical considerations: 1) 76% images we annotated are generated facial images, ensuring no potential for harm to any individual. 2) For real images, we only provide annotations for content either licensed by the original copyright holders or explicitly stated as freely shareable for research purposes.
Acknowledgments
This work is supported by the U.S. National Science Foundation (NSF) under grant IIS-2434967 and the National Artificial Intelligence Research Resource (NAIRR) Pilot and TACC Lonestar6. The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of NSF and NAIRR Pilot.
Appendix
Appendix A The Definition of Demographic Categories
Skin Tone: Skin tone is an important attribute of human appearance, with significant variation from pale to dark. Recently, AI systems, especially computer vision models, have become controversial over concerns about the potential bias of performance varying based on skin tone [102, 46, 103]. Additionally, existing research has pointed out that skin tone annotations can be potentially less biased than building a racial classifier [104]. And the ethnicity attribute is subjective and can conceptually cause confusion in many aspects; for example, there may be no difference in facial appearance of African-American and African people, although, they may be referred to with two distinct racial categories. We, therefore, have opted to annotate the apparent skin tone of each face image. The Monk Skin Tone Scale [48] is developed specifically for the computer vision use case. We intentionally use the Monk Skin Tone scale over the Fitzpatrick skin type [105], which is developed as means for determining one’s likelihood of getting sunburn and lacks variance in darker skin tones [106, 107]. Additionally, Fitzpatrick skin type has been shown to be unreliable for image annotation [108].
Gender: Many governments [49, 50] have adopted binary gender (i.e., Man/Male (M) and Woman/Female(F), defined as sex at birth, as a common choice for legal and institutional systems and official documents. Most facial recognition research [45, 51, 52] also considers binary genders in their analyses. Our AI-Face dataset adopts binary gender as gender attributes.
Age: Follow United Nations [53] and Statistics Canada [54], we have five distinct perceived age groups- Child (0-14), Youth (15-24), Adults (25-44), Middle-age Adults (45-64), and Seniors (65+).
The demographic attribute and its corresponding example images are shown from Fig. A.1 to Fig. A.3.
Figure A.1: Demographic annotation definition and examples of skin tone attribute.
Figure A.2: Demographic annotation definition and examples of gender attribute.
Figure A.3: Demographic annotation definition and examples of age attribute.
Appendix B The Details of Demographically Annotated AI-Face Dataset
Figure B.1: Overview of AI-Face dataset. Each face has three demographic annotations.
B.1 Detailed Information of Datasets
We build our AI-Face dataset by collecting and integrating public real and AI-generated face images sourced from academic publications, GitHub repositories, and commercial tools. We strictly adhere to the license agreements of all datasets to ensure that they allow inclusion in our datasets and secondary use for training and testing. Table B.1 shows the detailed information of each dataset we used in our AI-Face, including the number of samples, the link for downloading the dataset, the accessibility, and their licenses.
Dataset #Samples Link Access License
FF++
 [2]
127K https://github.com/ondyari/FaceForensics/tree/master/dataset
freely shared for a research purpose,
submit aggreement
Non Commercial
DFDC
 [21]
75K https://www.kaggle.com/c/deepfake-detection-challenge/data Unknown
DFD
 [22]
40K https://research.google/blog/contributing-data-to-deepfake-detection-research/ the rights have been cleared for real videos, submit aggreement Non Commercial
Celeb-DF-v2
 [23]
179K https://cse.buffalo.edu/\$\$siweilyu/celeb-deepfakeforensics.html
freely shared for a research purpose,
submit aggreement
Non Commercial
AttGAN
 [55]
6K
StarGAN
 [55]
5.6K
StyleGAN
 [55]
10K https://iplab.dmi.unict.it/mfs/Deepfakes/PaperGANDCT-2021/ Online dataset, download directly, no license or agreement to sign Unknown
StyleGAN2
 [57]
118K https://github.com/SelfishGene/SFHQ-dataset
Since all images in this dataset are synthetically
generated there are no privacy issues or
license issues surrounding these images.
MIT License
StyleGAN3
 [58]
26.7K https://huggingface.co/datasets/InfImagine/FakeImageDataset
This dataset are fully open for academic research
and can be used for commercial purposes
with official written permission.
Apache-2.0
MMDGAN
 [56]
1K
MSGGAN
 [56]
1K
STGAN
 [56]
1K https://github.com/vishal3477/Reverse_Engineering_GMs/blob/main/dataset/ The dataset can be used for research purposes only and can be used for commercial purposes with official written permission. Non Commercial
ProGAN
 [59]
100K https://drive.google.com/drive/folders/1jU-hzyvDZNn_M3ucuvs9xxtJNc9bPLGJ
Online dataset, download directly,
no license or agreement to sign
Unknown
VQGAN
 [60]
50K https://github.com/awsaf49/artifact
This dataset comes from ArtiFact dataset,
which dataset takes leverage of data
from multiple methods thus different parts
of the dataset come with different licenses.
MIT License
DALLE2
 [61]
204
IF
 [61]
505
Midjourney
 [61]
100 https://github.com/ZhendongWang6/DIRE freely shared for a research purpose
DCFACe
 [62]
529K https://github.com/mk-minchul/dcface freely shared for a research purpose Unknown
Latent Diffusion
 [63]
20K https://github.com/grip-unina/DMimageDetection
Copyright 2024 Image Processing Research Group of
University Federico II of Naples (’GRIP-UNINA’).
All rights reserved.Licensed under the
Apache License, Version 2.0 (the ”License”)
Apache-2.0
Palette
 [64]
6K https://github.com/awsaf49/artifact/?tab=readme-ov-file#data-generation
This dataset comes from ArtiFact dataset,
which dataset takes leverage of data from
multiple methods thus different parts
of the dataset come with different licenses.
MIT License
SD v1.5
 [65]
18K
SD Inpainting
 [65]
20.9K https://huggingface.co/datasets/OpenRL/DeepFakeFace freely shared for a research purpose Apache-2.0
FFHQ
 [6]
70K https://github.com/NVlabs/ffhq-dataset
You can use, redistribute, and adapt it for
non-commercial purposes, as long as you
(a) give appropriate credit by citing our paper,
(b) indicate any changes that you’ve made,and
(c) distribute any derivative works under the same license.
Creative Commons
BY-NC-SA 4.0 license
IMDB-WIKI
 [20]
239K https://data.vision.ee.ethz.ch/cvl/rrothe/imdb-wiki/
This dataset is made available for academic
research purpose only. All the images are
collected from the Internet, and the
copyright belongs to the original owners.
Non Commericial
Table B.1: A list of datasets used in AI-Face, including the number of samples, links, access details, and licenses.
B.2 Artifacts of Deepfake Forgeries in Frequency
Leveraging frequency domain information plays a pivotal role in detecting AI-generated images. Frequency-based methods analyze the frequency components of an image, capturing information that may not be readily apparent in the spatial domain. In Fig. B.2, we present the mean Fast Fourier Transform (FFT) spectrum of images sampled from various sources in our AI-Face dataset. The results indicate that generative models often concentrate their output energy in the low-frequency range, represented by the central area of the FFT spectrum, resulting in overly smooth images. Notably, some models, such as StarGAN and Midjourney, exhibit distinct frequency artifacts, suggesting that they continue to struggle with eliminating generative patterns in the frequency domain. These artifacts serve as critical cues for distinguishing synthetic images from real ones. While most prior work has focused on applying frequency information to enhance the utility performance of detectors, exploring how frequency features can be leveraged to improve the fairness of detectors presents a promising direction for future research.
Figure B.2: Frequency analysis on various sources. The mean FFT spectrum computation involves averaging over 2,000 images. DALLE2, IF, and Midjourney take average over 200, 500, and 100, respectively, due to their small number of samples.
B.3 Experimental Study of Existing Face Attribute Prediction Tools
We compare current state-of-the-art face attribute prediction tools Face++ [71] and InsightFace [72] with our annotator. We perform intra-domain (train and test on IMDB-WIKI) and cross-domain (train on IMDB-WIKI, test on four AI-generated face datasets) evaluations. FF++, DFDC, DFD, and Celeb-DF-v2 are selected for cross-domain evaluation because they contain AI-generated faces, which match our objective and are not used to train Face++ and InsightFace. Additionally, they have demographic attribute annotations from [19], which can be used as ground truth for annotator evaluation. Since those annotations provided by  [19] have limited age annotations, our evaluation of these four datasets is confined to gender. The intra-domain results are shown in Table B.2 and the results of cross-domain are in Table B.3. Those results demonstrate our annotator’s superiority in demographic attribute prediction and generalization capability against Face++ and InsightFace. For example, under intra-domain evaluation (Table B.2), its precision surpasses the second-best method, InsightFace, by 3.47% on Female and 24.81% on Senior. In cross-domain evaluation (Table B.2), our annotator maintains high accuracy on all datasets, reflecting good generalization. For instance, on the DFDC dataset, the precision our annotator outperforms Face++ by a margin of up to 1.07% and InsightFace by 3.32% on Female.
Female Male Child Young Adult Mid Senior
Method Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
0.9161 0.9142 0.9151 0.9143 0.9163 0.9153 0.8579 0.0220 0.0429 0.3942 0.2940 0.3368 0.3215 0.6700 0.4345 0.5423 0.6460 0.5894 0.8078 0.7700 0.7884
Face ++ (0.0064) (0.0031) 0.0045 (0.0033) (0.0067) 0.0048 (0.0962) (0.0062) (0.0119) (0.0222) (0.0161) (0.0186) (0.0054) (0.0192) (0.0087) (0.0160) (0.0210) (0.0147) (0.0143) (0.0165) (0.0127)
0.9648 0.9405 0.9525 0.9420 0.9656 0.9537 1.0000 0.0020 0.0040 0.3970 0.0693 0.1180 0.2599 0.6180 0.3659 0.4105 0.5733 0.4783 0.7224 0.7560 0.7386
Insightface (0.0055) (0.0058) 0.0017 (0.0050) (0.0057) 0.0017 (0.0000) (0.0016) (0.0032) (0.0337) (0.0098) (0.0155) (0.0027) (0.0124) (0.0048) (0.0067) (0.0196) (0.0094) (0.0124) (0.0238) (0.0143)
0.9995 0.9992 0.9993 0.9992 0.9995 0.9993 0.9787 0.9780 0.9783 0.9560 0.9393 0.9476 0.9265 0.9547 0.9402 0.9498 0.9320 0.9408 0.9642 0.9700 0.9671
Ours (0.0006) (0.0006) 0.0006 (0.0006) (0.0006) 0.0006 (0.0027) (0.0062) (0.0039) (0.0078) (0.0080) (0.0042) (0.0122) (0.0107) (0.0058) (0.0123) (0.0123) (0.0124) (0.0085) (0.0087) (0.0082)
Table B.2: Detailed comparison of our annotator against Face++ [71] and InsightFace [72] on IMDB-WIKI  [20] dataset. Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings are reported. The best results are shown in Bold.
Female Male
Dataset Method precision recall F1 precision recall F1
0.9816 0.9795 0.9805 0.9795 0.9816 0.9805
Face ++ (0.3021) (0.1360) (0.1459) (0.1312) (0.3084) (0.1508)
0.9700 0.9664 0.9682 0.9666 0.9713 0.9683
Insightface (0.4697) (0.6046) (0.3867) (0.5794) (0.4815) (0.3802)
0.9799 0.9992 0.9894 0.9992 0.9795 0.9892
FF++ Ours (0.0022) (0.0006) (0.0009) (0.0007) (0.0023) (0.0009)
0.9412 0.8992 0.9197 0.9035 0.9437 0.9231
Face ++ (0.9771) (1.0095) (0.5639) (0.8353) (1.0246) (0.5353)
0.9187 0.7869 0.8475 0.8139 0.9301 0.8680
Insightface (0.9855) (1.7976) (0.7444) (1.1401) (1.0587) (0.3842)
0.9519 0.9741 0.9629 0.9735 0.9507 0.0114
DFDC Ours (0.0106) (0.0014) (0.0059) (0.0015) (0.9619) (0.0064)
0.9501 0.8228 0.8818 0.8440 0.9568 0.8968
Face ++ (0.3773) (1.8758) (1.0113) (1.3784) (0.3907) (0.6856)
0.9441 0.7557 0.8394 0.7964 0.9552 0.8686
Insightface (0.7765) (0.9555) (0.7937) (0.6834) (0.6344) (0.5961)
0.9378 0.9365 0.9366 0.9366 0.9379 0.9372
DFD Ours (0.0045) (0.0053) (0.0033) (0.0049) (0.0048) (0.0033)
0.9989 0.9648 0.9815 0.9660 0.9989 0.9822
Face ++ (0.0553) (0.4182) (0.2361) (0.3918) (0.0533) (0.2215)
0.9984 0.9811 0.9896 0.9814 0.9984 0.9898
Insightface (0.0541) (0.3518) (0.1801) (0.3396) (0.0534) (0.1737)
1.0000 0.9997 0.9999 0.9997 1.0000 0.9999
Celeb-DF-v2 Ours (0.0000) (0.0005) (0.0003) (0.0005) (0.0000) (0.0003)
Table B.3: Detailed comparison of our annotator against Face++ [71] and InsightFace [72] on FF++ [2], DFDC [21], DFD [22], and Celeb-DF-v2 [23] datasets. Prediction mean and standard deviation (in parentheses) of each method across 5 random samplings. The best results are shown in Bold.
B.4 Annotator Implementation Detail
[…]
Figure B.3: Visualization of the skin tone estimation process.
B.5 Details of Human Labeling Activities in Annotation Quality Assessment
The annotation process for assessing the quality of AI-generated face image annotations followed a structured and ethically grounded methodology. Prior to labeling, all human annotators signed an Annotator Agreement outlining the project objectives, confidentiality requirements, and detailed labeling guidelines for gender and age classification. This agreement emphasized impartiality, respect, and adherence to professional conduct throughout the annotation activities. Human annotators then underwent tutorial training using real example images to familiarize themselves with demographic attributes and labeling criteria, focusing on identifying gender-specific features, such as facial structure and presence of facial hair, and age-related indicators, including wrinkles and skin elasticity.
Following the agreement and training, human annotators independently labeled the images based on the established criteria, categorizing gender and age into predefined classes, and recorded their annotations in a CSV file. A structured conflict resolution approach ensured accuracy and consistency in annotations. Labels agreed upon by a majority of annotators were finalized directly, while unanimous disagreements were resolved through collaborative discussions guided by the annotation guidelines. This process ensured that all annotations were objective, reliable, and aligned with ethical standards set forth in the signed agreement.
B.6 Visualization of Skin Tone Annotation Generation
The visualization shown in Fig. B.3 illustrates the skin tone estimation process using the Monk Skin Tone (MST) Scale. Each row represents a sample image, showing the progression from the original face with facial landmarks to the masked skin region that excludes non-skin areas like eyes and lips. Subsequently, the K-Means clustered skin region highlights the dominant skin tones extracted from the facial area. On the right, bar plots display the proportions of the top three dominant tones within the clustered region, with the top tone (largest cluster) mapped to the closest MST Scale shade. This mapping is achieved by calculating the maximum similarity, as indicated by the Euclidean distance in RGB space between the cluster centroid and MST reference colors. This process visually demonstrates how the methodology isolates, clusters, and estimates skin tones for accurate skin tone annotation generation.
Appendix C Fairness Benchmark Settings
C.1 Implementation Detail
For fairness benchmark, all experiments are based on the PyTorch with a single NVIDIA RTX A6000 GPU. During training, we utilize SGD optimizer with a learning rate of 0.0005, with momentum of 0.9 and weight decay of 0.005. The batch size is set to 128 for most detectors. However, for the SRM [87], UCF [26], and PG-FDD [30], the batch size is adjusted to 32 due to GPU memory. For hyperparameters defined in these detectors, we use the default values set in their original papers. All detectors are initialized with their official pre-trained weights, and trained for 10 epochs.
C.2 Details of Detection Methods
We summarized the backbone architecture, GitHub repository link, and publication venue of the detectors implemented in our fairness benchmark in Table C.1. A brief introduction to each detector is provided below:
Model Type Detector Backbone GitHub Link VENUE
Naive Xception [82] Xception https://github.com/ondyari/FaceForensics/blob/master ICCV-2019
Efficient-B4 [83] EfficientNet https://github.com/lukemelas/EfficientNet-PyTorch ICML-2019
ViT-B/16 [84] Transformer https://github.com/lucidrains/vit-pytorch ICLR-2021
Spatial UCF [26] Xception https://github.com/SCLBD/DeepfakeBench/tree/main ICCV-2023
UnivFD [88] CLIP VIT https://github.com/Yuheng-Li/UniversalFakeDetect CVPR-2023
CORE [89] Xception https://github.com/niyunsheng/CORE CVPRW-2022
Frequency F3Net [85] Xception https://github.com/yyk-wew/F3Net ECCV-2020
SRM [87] Xception https://github.com/SCLBD/DeepfakeBench/tree/main CVPR-2021
SPSL [86] Xception https://github.com/SCLBD/DeepfakeBench/tree/main CVPR-2021
Fairness- enhanced DAW-FDD [29] Xception Unpublished code, reproduced by us WACV-2024
DAG-FDD [29] Xception Unpublished code, reproduced by us WACV-2024
PG-FDD [30] Xception https://github.com/Purdue-M2/Fairness-Generalization CVPR-2024
Table C.1: Summary of the implemented detectors in our fairness benchmark.
Xception [82]: is a deep convolutional neural network (CNN) architecture that relies on depthwise separable convolutions. This approach significantly reduces the number of parameters and computational cost while maintaining high performance. Xception serves as a classic backbone in deepfake detectors.
EfficientB4 [83]: is part of the EfficientNet family [83], which utilizes a novel model scaling method that uniformly scales all dimensions of depth, width, and resolution using a compound coefficient. EfficientNet also serves as a classic backbone in deepfake detectors.
ViT-B/16 [84]: is a model that applies the transformer architecture, the ’B’ denotes the base model size, and ’16’ indicates the patch size. ViT-B/16 splits images into 16 patches, linearly embeds each patch, adds positional embeddings, and feeds the resulting sequence of vectors into a standard transformer encoder.
F3Net [85]: utilizes a cross-attention two-stream network to effectively identify frequency-aware clues by integrating two branches: FAD and LFS. The FAD (Frequency-aware Decomposition) module divides the input image into various frequency bands using learnable partitions, representing the image with frequency-aware components to detect forgery patterns through this decomposition. Meanwhile, the LFS (Localized Frequency Statistics) module captures local frequency statistics to highlight statistical differences between authentic and counterfeit faces.
SPSL [86]: integrates spatial image data with the phase spectrum to detect up-sampling artifacts in face forgeries, enhancing the model’s generalization ability for face forgery detection. The paper provides a theoretical analysis of the effectiveness of using the phase spectrum. Additionally, it highlights that local texture information is more important than high-level semantic information for accurately detecting face forgeries.
SRM [87]: extracts high-frequency noise features and combines two different representations from the RGB and frequency domains to enhance the model’s generalization ability for face forgery detection.
UCF [26]: presents a multi-task disentanglement framework designed to tackle two key challenges in deepfake detection: overfitting to irrelevant features and overfitting to method-specific textures. By identifying and leveraging common features, this framework aims to improve the model’s generalization ability.
UnivFD [88]: uses the frozen CLIP ViT-L/14 [73] as feature extractor and trains the last linear layer to classify fake and real images.
CORE [89]: explicitly enforces the consistency of different representations. It first captures various representations through different augmentations and then regularizes the cosine distance between these representations to enhance their consistency.
DAW-FDD [29]: a demographic-aware Fair Deepfake Detection (DAW-FDD) method leverages demographic information and employs an existing fairness risk measure [109]. At a high level, DAW-FDD aims to ensure that the losses achieved by different user-specified groups of interest (e.g., different races or genders) are similar to each other (so that the AI face detector is not more accurate on one group vs another) and, moreover, that the losses across all groups are low. Specifically, DAW-FDD uses a CVaR [110, 111] loss function across groups (to address imbalance in demographic groups) and, per group, DAW-FDD uses another CVaR loss function (to address imbalance in real vs AI-generated training examples).
DAG-FDD [29]: a demographic-agnostic Fair Deepfake Detection (DAG-FDD) method, which is based on the distributionally robust optimization (DRO) [112, 113]. To use DAG-FDD, the user does not have to specify which attributes to treat as sensitive such as race and gender, only need to specify a probability threshold for a minority group without explicitly identifying all possible groups.
PG-FDD [30]: PG-FDD (Preserving Generalization Fair Deepfake Detection) employs disentanglement learning to extract demographic and domain-agnostic forgery features, promoting fair learning across a flattened loss landscape. Its framework combines disentanglement learning, fairness learning, and optimization modules. The disentanglement module introduces a loss to expose demographic and domain-agnostic features that enhance fairness generalization. The fairness learning module combines these features to promote fair learning, guided by generalization principles. The optimization module flattens the loss landscape, helping the model escape suboptimal solutions and strengthen fairness generalization.
C.3 Fairness Metrics
We assume a test set comprising indices {1, …, $n$}. $Y_{j}$ and $\hat{Y}_{j}$ respectively represent the true and predicted labels of the sample $X_{j}$. Their values are binary, where 0 means real and 1 means fake. For all fairness metrics, a lower value means better performance. The formulations of fairness metrics are as follows,
$F_{EO}:=\sum_{\mathcal{J}_{j}\in\mathcal{J}}\sum_{q=0}^{1}\left|\frac{\sum_{j=1}^{n}\mathbb{I}_{\left[\hat{Y}_{j}=1,D_{j}=\mathcal{J}_{j},Y_{j}=q\right]}}{\sum_{j=1}^{n}\mathbb{I}_{\left[D_{j}=\mathcal{J}_{j},Y_{j}=q\right]}}-\frac{\sum_{j=1}^{n}\mathbb{I}_{\left[\hat{Y}_{j}=1,Y_{j}=q\right]}}{\sum_{j=1}^{n}\mathbb{I}_{\left[Y_{j}=q\right]}}\right|,$
$F_{O\!A\!E}:=\max_{\mathcal{J}_{j}\in\mathcal{J}}\left\{\frac{\sum_{j=1}^{n}\mathbb{I}_{[\hat{Y}_{j}=Y_{j},D_{j}=\mathcal{J}_{j}]}}{\sum_{j=1}^{n}\mathbb{I}_{[D_{j}=\mathcal{J}_{j}]}}\right.\quad\left.-\min_{{\mathcal{J}_{j}}^{\prime}\in\mathcal{J}}\frac{\sum_{j=1}^{n}\mathbb{I}_{[\hat{Y}_{j}=Y_{j},D_{j}={\mathcal{J}_{j}}^{\prime}]}}{\sum_{j=1}^{n}\mathbb{I}_{[D_{j}={\mathcal{J}_{j}}^{\prime}]}}\right\},$
$F_{DP}:=\max_{q\in\{0,1\}}\left\{\max_{J_{j}\in\mathcal{J}}\frac{\sum_{j=1}^{n}\mathbb{I}_{[\hat{Y}_{j}=q,D_{j}=J_{j}]}}{\sum_{j=1}^{n}\mathbb{I}_{[D_{j}=J_{j}]}}\right.\quad\left.-\min_{J_{j}^{\prime}\in\mathcal{J}}\frac{\sum_{j=1}^{n}\mathbb{I}_{[\hat{Y}_{j}=q,D_{j}=J_{j}^{\prime}]}}{\sum_{j=1}^{n}\mathbb{I}_{[D_{j}=J_{j}^{\prime}]}}\right\},$
$F_{M\!E\!O}:=\max_{q,q^{\prime}\in\{0,1\}}\left\{\max_{J_{j}\in\mathcal{J}}\frac{\sum_{j=1}^{n}\mathbb{I}_{[\hat{Y}_{j}=q,Y_{j}=q^{\prime},D_{j}=J_{j}]}}{\sum_{j=1}^{n}\mathbb{I}_{[D_{j}=J_{j},Y_{j}=q]}}\right.\quad\left.-\min_{J_{j}^{\prime}\in\mathcal{J}}\frac{\sum_{j=1}^{n}\mathbb{I}_{[\hat{Y}_{j}=q,Y_{j}=q^{\prime},D_{j}=J_{j}^{\prime}]}}{\sum_{j=1}^{n}\mathbb{I}_{[D_{j}=J_{j}^{\prime},Y_{j}=q]}}\right\},$
$F_{IND}:=\sum_{j=1}^{n-1}\sum_{l=j+1}^{n}[\left|f(X_{j})-f(X_{l})\right|-\delta\|X_{j}-X_{l}\|_{2}]_{+},$
where $D$ is the demographic variable, $\mathcal{J}$ is the set of subgroups with each subgroup $\mathcal{J}_{j}\in\mathcal{J}$. $M$ is the set of detection models and $F$ is the set of fairness metrics. $F_{EO}$ measures the disparity in TPR and FPR between each subgroup and the overall population. $F_{OAE}$ measures the maximum ACC gap across all demographic groups. $F_{DP}$ measures the maximum difference in prediction rates across all demographic groups. And $F_{MEO}$ captures the largest disparity in prediction outcomes (either positive or negative) when comparing different demographic groups. $\delta$ in $F_{IND}$ is a predefined scale factor (0.08 in our experiments). $[\cdot]_{+}$ is the hinge function, $\|\cdot\|_{2}$ is the $\ell_{2}$ norm. $f(X_{j})$ represents the predicted logits of the model for input sample $X_{j}$. $F_{IND}$ points that a model should be fair across individuals if similar individuals have similar predicted outcomes.
Appendix D More Fairness Benchmark Results and Analysis
D.1 Detailed Results of Overall Performance Comparison
Detailed test results of each subgroup of each detector on AI-Face are presented in this section. Table D.1 provides comprehensive metrics of each subgroup on AI-Face. These results and findings align with the results reported in Table. 4 submitted main manuscript.
Model Type Method Metric Gender Skin Tone Age Intersection
M F L M D Child Young Adult Middle Senior M-L M-M M-D F-L F-M F-D
Naive Xception  [82] AUC 98.90 98.20 97.69 98.44 98.88 95.95 97.86 99.10 98.66 96.54 97.88 98.70 99.22 97.53 98.17 98.19
FPR 11.12 15.10 18.37 14.72 9.54 36.83 18.26 8.95 12.78 20.42 17.70 12.82 8.42 18.93 16.58 11.32
TPR 99.38 99.22 98.94 99.62 98.43 99.55 99.38 99.21 99.49 98.51 98.55 99.64 98.87 99.16 99.59 97.58
ACC 96.77 95.80 95.16 96.43 96.03 89.83 94.98 97.10 97.06 92.06 94.26 96.79 96.78 95.70 96.10 94.67
EfficientB4  [83] AUC 98.86 98.31 99.23 98.94 97.59 99.63 98.61 98.44 98.82 98.69 99.01 99.10 98.36 99.39 98.78 95.94
FPR 17.57 22.96 21.32 17.17 25.47 16.13 22.92 19.79 19.41 19.07 20.50 15.44 20.74 22.00 18.87 33.02
TPR 99.04 98.56 99.49 99.02 98.19 99.56 98.89 98.55 99.14 98.66 99.22 99.25 98.60 99.64 98.81 97.40
ACC 94.91 93.42 94.95 95.43 91.05 95.37 93.45 93.81 95.48 92.62 94.01 95.90 93.04 95.51 94.99 87.44
ViT-B/16  [84] AUC 99.02 98.26 97.50 98.77 98.49 95.55 98.23 98.98 98.80 97.28 97.21 99.03 99.03 97.80 98.50 97.21
FPR 14.06 19.17 21.74 16.86 15.48 28.28 21.09 13.34 17.22 20.92 21.43 14.93 12.61 22.00 18.76 20.05
TPR 98.43 97.51 96.75 98.23 97.36 94.58 97.88 98.14 98.40 96.16 96.77 98.56 98.21 96.74 97.93 95.72
ACC 95.33 93.52 92.72 94.88 93.49 88.48 93.16 95.17 95.31 90.34 91.96 95.48 95.10 93.16 94.33 90.55
Frequency F3Net  [85] AUC 99.09 98.24 98.51 98.68 98.79 96.05 98.17 99.15 98.97 97.40 98.66 99.00 99.27 98.30 98.36 97.80
FPR 12.49 17.21 30.72 16.54 10.78 49.32 21.21 10.51 12.52 20.00 30.75 14.33 9.38 30.69 18.72 13.02
TPR 99.15 99.00 99.69 99.41 98.11 99.77 99.40 99.00 99.06 97.96 99.55 99.34 98.76 99.76 99.48 96.86
ACC 96.25 95.12 93.05 95.87 95.43 86.66 94.27 96.54 96.77 91.84 91.55 96.22 96.42 93.95 95.55 93.63
SPSL  [86] AUC 98.88 98.58 98.99 98.70 98.81 97.41 98.05 99.17 98.75 97.32 99.05 98.78 99.02 98.92 98.61 98.36
FPR 11.62 16.03 19.50 14.77 11.44 37.20 20.20 9.58 13.12 19.07 18.94 12.88 9.57 19.95 16.62 14.43
TPR 99.64 99.52 99.53 99.73 99.15 99.78 99.58 99.53 99.69 99.12 99.44 99.77 99.37 99.58 99.70 98.73
ACC 96.84 95.80 95.38 96.51 95.96 89.90 94.65 97.17 97.16 92.92 94.59 96.88 96.80 95.85 96.17 94.42
SRM  [87] AUC 98.45 97.40 98.71 97.94 97.95 97.24 97.27 98.42 98.13 97.78 99.24 98.46 98.52 98.36 97.48 96.82
FPR 12.84 19.11 22.30 17.50 12.30 36.76 22.42 11.92 15.33 19.64 18.94 14.70 9.92 25.06 20.25 16.10
TPR 98.84 98.33 99.41 99.02 97.35 99.23 98.91 98.29 98.97 97.52 99.89 99.20 98.06 99.16 98.85 95.99
ACC 95.93 94.16 94.67 95.35 94.44 89.62 93.59 95.65 96.14 91.67 94.91 96.03 95.76 94.53 94.72 92.03
Spatial UCF  [26] AUC 98.62 97.45 97.20 97.92 98.49 95.59 97.26 98.74 98.67 97.04 97.67 98.44 99.00 96.88 97.41 97.47
FPR 11.30 16.37 26.23 16.01 8.90 55.93 21.10 8.43 11.70 20.40 24.85 13.82 7.23 27.37 18.16 11.57
TPR 98.20 97.77 98.75 98.37 96.89 99.53 98.50 97.58 98.35 96.91 99.00 98.47 97.63 98.61 98.28 95.45
ACC 95.84 94.39 93.30 95.18 95.14 84.71 93.61 96.03 96.36 91.01 92.70 95.66 96.24 93.65 94.73 93.15
UnivFD  [88] AUC 98.55 97.76 98.38 98.47 97.40 98.13 98.07 98.33 98.11 96.94 98.24 98.76 98.13 98.46 98.19 95.84
FPR 16.46 20.97 20.06 19.04 17.60 19.90 19.91 16.70 22.13 17.04 17.70 16.84 15.88 22.00 21.20 20.36
TPR 98.02 97.12 97.57 98.26 95.69 97.43 97.40 97.39 98.41 94.49 97.10 98.54 96.99 97.83 98.02 93.17
ACC 94.42 92.80 93.73 94.43 91.68 92.80 93.09 93.74 94.35 90.56 93.19 95.03 93.29 94.04 93.87 88.74
CORE  [89] AUC 99.04 98.01 97.80 98.47 98.79 95.88 97.82 99.09 98.77 97.11 98.12 98.91 99.26 97.57 98.02 97.84
FPR 10.73 16.52 21.46 14.76 10.68 43.07 19.42 9.19 12.34 20.10 18.63 12.14 8.45 23.79 17.34 14.25
TPR 99.40 99.27 99.61 99.51 98.84 99.83 99.27 99.27 99.46 99.00 99.78 99.53 99.13 99.52 99.49 98.28
ACC 96.88 95.49 95.01 96.34 95.97 88.37 94.61 97.08 97.13 92.49 94.91 96.87 96.95 95.07 95.85 94.17
Fairness- enhanced DAW-FDD  [29] AUC 98.36 97.15 96.84 97.45 98.51 94.97 96.41 98.62 98.00 94.95 96.83 98.05 98.85 96.77 96.87 97.91
FPR 14.12 19.63 26.93 18.59 12.81 56.69 23.64 11.23 15.40 26.17 26.40 16.16 10.80 27.37 20.98 16.03
TPR 99.42 99.24 99.73 99.38 99.20 99.77 99.18 99.28 99.51 98.98 99.78 99.47 99.32 99.70 99.29 98.97
ACC 96.05 94.73 93.91 95.39 95.58 84.69 93.49 96.56 96.56 90.41 92.86 95.90 96.41 94.53 94.91 94.06
DAG-FDD  [29] AUC 99.05 98.44 97.56 98.79 98.73 96.55 98.14 99.10 98.91 97.59 98.11 99.00 99.16 97.10 98.57 97.83
FPR 12.11 18.02 20.76 15.10 14.21 26.95 20.01 11.72 14.61 22.40 18.32 13.04 10.58 22.76 17.14 20.00
TPR 99.21 99.05 98.98 99.25 98.83 97.64 98.92 99.18 99.37 99.13 98.77 99.33 98.98 99.10 99.17 98.54
ACC 96.39 94.97 94.67 96.06 94.90 91.07 94.20 96.36 96.61 91.79 94.26 96.51 96.23 94.92 95.65 92.47
PG-FDD  [30] AUC 99.36 98.94 98.94 99.18 99.13 98.83 98.89 99.35 99.27 97.85 98.97 99.35 99.39 98.89 99.02 98.51
FPR 9.49 12.68 15.29 12.06 8.82 22.77 14.31 7.97 11.64 19.74 13.35 10.90 7.30 16.88 13.20 11.26
TPR 98.73 98.21 98.63 98.85 97.43 99.11 98.44 98.28 98.87 97.74 98.66 99.02 98.12 98.61 98.69 96.10
ACC 96.68 95.61 95.59 96.43 95.55 93.27 95.26 96.66 96.79 91.78 95.49 96.75 96.56 95.65 96.12 93.69
Table D.1: Detailed test results of each subgroup of each detector on the AI-Face. In the Skin Tone groups, ‘L’ represents Light (Tone 1-3), ‘M’ is Medium (4-6), ‘D’ is Dark (Tone 7-10).
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 10.901 4.384 17.219 14.583 9.620 15.508 14.978 2.441 13.135 12.519 12.597 13.965
$F_{DP}$ 11.274 9.191 10.713 12.739 11.711 11.282 11.549 8.117 11.968 11.179 12.026 11.768
$F_{OAE}$ 2.434 3.609 2.276 2.232 2.814 1.780 1.950 2.940 1.658 0.878 1.753 1.439
$F_{EO}$ 0.160 0.093 0.205 0.209 0.156 0.191 0.186 0.034 0.176 0.165 0.176 0.176
Gender $F_{MEO}$ 5.475 5.458 8.003 5.749 5.754 5.848 5.575 3.244 4.367 5.186 5.808 4.086
$F_{DP}$ 1.205 1.412 2.416 1.340 1.445 1.959 1.810 0.781 0.980 1.715 1.470 1.545
$F_{OAE}$ 2.043 1.800 1.896 2.054 1.969 1.471 1.569 1.413 1.848 1.430 2.012 1.133
$F_{EO}$ 0.066 0.063 0.083 0.068 0.067 0.062 0.060 0.043 0.056 0.056 0.068 0.044
Age $F_{MEO}$ 28.244 7.460 38.521 27.860 24.768 40.542 44.342 8.584 34.156 36.450 35.031 36.197
$F_{DP}$ 11.228 12.245 12.140 11.395 11.466 14.564 15.856 16.134 13.525 12.256 13.478 12.082
$F_{OAE}$ 7.138 5.234 10.940 6.933 6.053 11.126 11.481 4.171 8.294 9.192 8.636 8.934
$F_{EO}$ 0.460 0.175 0.560 0.460 0.410 0.550 0.610 0.191 0.508 0.537 0.524 0.539
Intersection $F_{MEO}$ 15.752 10.644 24.460 18.455 15.157 18.381 17.397 5.300 16.257 14.806 15.219 16.517
$F_{DP}$ 16.943 14.565 13.773 18.071 17.490 14.943 15.612 12.967 17.063 14.802 16.786 15.513
$F_{OAE}$ 6.805 8.029 5.025 6.658 7.200 3.614 3.532 6.226 5.079 3.314 5.757 2.989
$F_{EO}$ 0.355 0.307 0.441 0.440 0.366 0.382 0.382 0.178 0.371 0.336 0.399 0.354
Utility(%) - AUC 0.968 0.968 0.981 0.966 0.968 0.967 0.977 0.979 0.975 0.970 0.973 0.978
ACC 0.931 0.922 0.924 0.930 0.929 0.946 0.951 0.933 0.947 0.941 0.941 0.952
AP 0.987 0.988 0.994 0.986 0.988 0.985 0.991 0.993 0.989 0.987 0.989 0.991
EER 0.093 0.101 0.082 0.096 0.095 0.076 0.074 0.083 0.076 0.085 0.084 0.072
FPR 0.205 0.219 0.290 0.199 0.190 0.205 0.163 0.188 0.144 0.186 0.168 0.151
Table D.2: Detailed fairness and utility evaluation results on 20% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 9.815 4.414 12.194 10.801 9.275 24.037 12.299 2.568 15.734 12.628 10.463 10.982
$F_{DP}$ 10.080 10.413 9.475 11.137 10.644 12.550 10.632 8.322 11.251 10.157 10.456 10.928
$F_{OAE}$ 0.122 0.095 0.146 0.154 0.122 0.281 0.150 0.043 0.180 0.154 0.135 0.143
$F_{EO}$ 1.472 3.796 3.395 2.088 1.631 3.280 1.533 2.898 2.384 1.571 1.323 1.188
Gender $F_{DP}$ 5.576 3.592 6.089 5.985 4.368 7.959 4.400 3.438 5.234 5.960 4.797 5.390
$F_{MEO}$ 1.817 0.822 1.853 1.566 1.227 2.658 1.481 0.797 1.866 2.053 1.458 2.052
$F_{OAE}$ 1.559 1.622 1.829 1.966 1.595 1.722 1.338 1.507 1.369 1.473 1.530 1.303
$F_{EO}$ 0.060 0.047 0.067 0.069 0.052 0.080 0.049 0.045 0.055 0.062 0.055 0.055
Age $F_{MEO}$ 32.781 9.931 18.050 26.665 33.004 54.967 43.829 7.840 38.202 41.707 34.285 33.582
$F_{DP}$ 12.161 12.428 14.762 10.954 13.272 16.006 14.243 16.076 13.394 13.661 11.630 12.955
$F_{OAE}$ 8.535 4.312 6.949 6.831 8.210 14.795 11.248 4.185 9.906 10.841 8.738 8.520
$F_{EO}$ 0.474 0.235 0.348 0.432 0.470 0.738 0.611 0.179 0.541 0.585 0.509 0.500
Intersection $F_{MEO}$ 13.451 8.558 15.615 13.559 12.356 30.133 15.278 5.585 19.342 17.006 12.784 13.197
$F_{DP}$ 13.795 16.478 14.409 16.096 15.138 15.893 14.797 13.263 15.123 14.586 14.345 15.166
$F_{OAE}$ 4.424 7.462 5.221 6.173 4.728 5.030 2.886 6.298 3.741 3.541 4.402 2.911
$F_{EO}$ 0.312 0.236 0.416 0.360 0.309 0.550 0.304 0.185 0.374 0.324 0.310 0.301
Utility(%) - AUC 0.979 0.976 0.980 0.974 0.982 0.957 0.979 0.979 0.981 0.976 0.978 0.982
ACC 0.951 0.940 0.933 0.937 0.951 0.938 0.955 0.934 0.957 0.948 0.949 0.960
AP 0.991 0.991 0.993 0.989 0.993 0.983 0.992 0.993 0.991 0.989 0.991 0.992
EER 0.068 0.078 0.081 0.083 0.066 0.117 0.072 0.082 0.058 0.074 0.074 0.055
FPR 0.165 0.147 0.136 0.186 0.147 0.245 0.157 0.180 0.151 0.184 0.169 0.130
Table D.3: Detailed fairness and utility evaluation results on 40% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 9.086 14.704 4.388 15.303 6.813 14.516 14.952 2.186 9.689 13.488 9.672 4.108
$F_{DP}$ 10.232 11.784 7.714 11.225 9.979 14.909 13.116 8.004 10.666 10.844 10.054 8.575
$F_{OAE}$ 1.531 2.017 2.572 2.320 1.247 1.733 1.562 2.777 1.208 1.664 1.259 1.383
$F_{EO}$ 0.124 0.194 0.084 0.177 0.100 0.234 0.208 0.043 0.131 0.163 0.125 0.055
Gender $F_{MEO}$ 4.418 6.743 8.445 5.545 5.242 7.331 5.713 4.182 4.395 5.579 4.978 2.430
$F_{DP}$ 1.243 2.063 2.697 2.096 1.736 3.656 1.845 1.142 1.622 2.153 1.600 1.061
$F_{OAE}$ 1.567 1.846 2.052 1.318 1.499 0.630 1.651 1.489 1.217 1.267 1.470 0.858
$F_{EO}$ 0.053 0.072 0.086 0.057 0.057 0.087 0.062 0.050 0.048 0.056 0.055 0.029
Age $F_{MEO}$ 35.231 27.998 17.573 42.366 35.428 37.043 34.243 5.520 28.666 40.326 38.409 24.355
$F_{DP}$ 12.874 12.663 11.691 15.070 12.924 16.570 13.579 15.256 11.264 13.929 13.112 10.449
$F_{OAE}$ 8.954 7.379 7.004 10.921 8.915 8.078 8.232 3.900 7.519 10.379 9.909 6.870
$F_{EO}$ 0.514 0.411 0.320 0.585 0.519 0.528 0.520 0.134 0.438 0.570 0.558 0.362
Intersection $F_{MEO}$ 11.554 18.923 10.063 18.907 10.175 20.404 18.818 5.414 12.995 15.944 12.552 5.425
$F_{DP}$ 14.625 15.884 9.908 15.240 14.093 18.967 18.087 12.890 15.331 14.584 13.949 13.379
$F_{OAE}$ 4.755 4.997 5.459 3.617 4.210 2.965 4.832 6.152 3.747 3.106 4.102 3.306
$F_{EO}$ 0.279 0.405 0.311 0.362 0.251 0.477 0.431 0.200 0.276 0.331 0.282 0.159
Utility(%) - AUC 0.976 0.980 0.986 0.981 0.983 0.976 0.981 0.981 0.982 0.977 0.982 0.982
ACC 0.948 0.945 0.943 0.961 0.952 0.927 0.951 0.935 0.956 0.960 0.950 0.960
AP 0.989 0.992 0.996 0.991 0.993 0.991 0.992 0.994 0.992 0.987 0.993 0.991
EER 0.074 0.071 0.065 0.058 0.067 0.092 0.068 0.078 0.062 0.060 0.069 0.058
FPR 0.166 0.183 0.177 0.137 0.154 0.137 0.137 0.198 0.145 0.143 0.172 0.127
Table D.4: Detailed fairness and utility evaluation results on 60% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 15.463 6.826 7.442 13.642 4.221 9.425 13.574 2.368 13.487 10.127 8.763 7.613
$F_{DP}$ 11.994 10.440 8.378 11.211 8.998 6.875 11.085 8.231 11.506 10.182 8.920 11.914
$F_{OAE}$ 1.998 2.390 1.540 1.713 2.141 7.511 1.661 2.777 1.538 1.375 1.484 1.822
$F_{EO}$ 0.192 0.107 0.095 0.171 0.064 0.187 0.168 0.036 0.170 0.119 0.114 0.129
Gender $F_{MEO}$ 4.209 3.639 9.461 5.189 4.116 2.328 5.402 4.084 5.058 4.112 4.143 4.035
$F_{DP}$ 1.171 1.043 3.025 1.749 1.191 2.803 1.778 1.082 1.899 1.537 1.560 1.401
$F_{OAE}$ 1.579 1.395 2.163 1.499 1.507 1.248 1.560 1.519 1.277 1.218 1.158 1.353
$F_{EO}$ 0.051 0.045 0.095 0.057 0.050 0.037 0.059 0.050 0.053 0.045 0.045 0.046
Age $F_{MEO}$ 33.930 16.272 10.222 45.076 20.048 11.857 45.508 5.788 37.055 30.409 29.707 20.058
$F_{DP}$ 15.167 11.254 11.643 15.938 11.379 8.360 15.357 15.620 13.565 11.925 9.880 10.487
$F_{OAE}$ 8.520 4.738 5.662 11.409 5.123 10.447 11.546 3.777 9.454 7.784 8.513 4.450
$F_{EO}$ 0.488 0.286 0.231 0.623 0.322 0.228 0.625 0.136 0.526 0.441 0.459 0.357
Intersection $F_{MEO}$ 19.488 11.396 16.807 15.829 6.631 12.599 16.926 5.116 16.594 13.597 11.226 10.523
$F_{DP}$ 17.093 15.841 11.946 15.708 13.376 11.010 15.607 13.121 15.834 14.758 12.474 16.906
$F_{OAE}$ 3.583 5.700 5.469 3.139 5.452 10.020 2.878 6.209 2.852 2.313 2.755 4.348
$F_{EO}$ 0.392 0.262 0.340 0.349 0.225 0.420 0.342 0.188 0.345 0.254 0.226 0.294
Utility(%) - AUC 0.985 0.985 0.987 0.981 0.984 0.979 0.982 0.981 0.984 0.981 0.987 0.988
ACC 0.950 0.949 0.940 0.958 0.950 0.816 0.956 0.936 0.959 0.963 0.960 0.953
AP 0.994 0.994 0.996 0.992 0.994 0.992 0.993 0.994 0.993 0.989 0.995 0.996
EER 0.065 0.064 0.062 0.069 0.064 0.066 0.068 0.078 0.060 0.053 0.057 0.058
FPR 0.145 0.162 0.206 0.134 0.148 0.039 0.141 0.189 0.138 0.118 0.144 0.099
Table D.5: Detailed fairness and utility evaluation results on 80% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 9.750 11.908 5.095 12.896 5.859 18.083 10.363 5.926 10.955 9.570 8.980 3.889
$F_{DP}$ 16.824 14.637 12.547 17.621 14.479 19.676 15.803 8.744 16.109 16.063 15.700 12.899
$F_{OAE}$ 2.401 2.973 1.714 3.630 0.905 5.752 3.249 5.302 3.524 2.641 2.684 1.164
$F_{EO}$ 0.164 0.155 0.119 0.183 0.113 0.248 0.147 0.083 0.149 0.142 0.132 0.069
Gender $F_{MEO}$ 3.535 0.952 4.275 4.788 4.439 4.201 4.655 6.898 3.981 3.848 3.674 3.792
$F_{DP}$ 2.241 0.715 3.129 3.184 3.060 2.245 3.426 4.328 2.850 2.562 2.514 2.790
$F_{OAE}$ 2.482 0.833 2.263 2.690 2.496 2.920 2.332 3.028 2.236 2.451 2.360 2.166
$F_{EO}$ 0.051 0.011 0.047 0.057 0.052 0.062 0.049 0.070 0.048 0.051 0.049 0.045
Age $F_{MEO}$ 23.282 14.195 13.627 32.679 26.796 52.939 30.332 8.251 32.612 23.934 23.834 15.408
$F_{DP}$ 17.521 18.304 16.426 20.437 18.489 28.303 18.892 12.343 19.913 17.103 18.221 15.255
$F_{OAE}$ 11.063 3.474 9.185 16.411 13.102 26.805 15.234 6.602 16.794 11.909 11.523 7.893
$F_{EO}$ 0.381 0.203 0.303 0.490 0.436 0.769 0.472 0.162 0.469 0.378 0.370 0.273
Intersection $F_{MEO}$ 11.921 17.564 7.727 17.622 10.201 21.679 15.031 12.980 12.775 13.010 11.104 7.173
$F_{DP}$ 23.484 21.311 17.501 23.519 21.115 24.119 22.167 13.434 21.787 22.542 22.120 18.247
$F_{OAE}$ 4.000 4.650 3.969 5.957 4.306 9.974 5.003 11.530 5.750 3.903 4.600 4.426
$F_{EO}$ 0.344 0.335 0.304 0.378 0.250 0.515 0.324 0.270 0.325 0.290 0.303 0.196
Utility(%) - AUC 0.984 0.984 0.984 0.982 0.982 0.935 0.981 0.979 0.986 0.983 0.987 0.989
ACC 0.937 0.892 0.922 0.928 0.929 0.900 0.930 0.830 0.933 0.937 0.944 0.940
AP 0.980 0.986 0.986 0.977 0.978 0.910 0.977 0.980 0.982 0.980 0.984 0.987
EER 0.062 0.064 0.067 0.061 0.066 0.129 0.065 0.082 0.056 0.061 0.053 0.052
FPR 0.087 0.005 0.111 0.122 0.111 0.164 0.116 0.337 0.115 0.095 0.083 0.095
Table D.6: Detailed fairness and utility evaluation results on a training subset with the ratio of real vs fake is 1:1.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 11.678 10.565 8.595 12.629 11.790 17.068 9.661 5.615 11.138 11.726 8.680 6.435
$F_{DP}$ 14.133 12.859 10.579 15.157 13.983 16.962 14.081 8.438 14.161 14.724 13.388 13.256
$F_{OAE}$ 4.539 3.671 4.407 4.104 4.894 5.036 3.931 5.461 4.389 3.379 3.006 2.232
$F_{EO}$ 0.128 0.151 0.102 0.148 0.127 0.200 0.107 0.081 0.121 0.141 0.103 0.081
Gender $F_{MEO}$ 6.942 2.295 10.586 7.818 7.327 8.990 6.054 6.525 7.518 6.259 5.934 4.944
$F_{DP}$ 4.378 0.094 6.203 4.881 4.632 5.572 4.086 4.136 4.867 4.093 3.917 3.565
$F_{OAE}$ 3.225 1.508 4.799 3.669 3.303 4.060 2.906 2.842 3.445 2.890 2.795 2.407
$F_{EO}$ 0.072 0.023 0.106 0.080 0.074 0.091 0.063 0.067 0.075 0.065 0.062 0.052
Age $F_{MEO}$ 36.384 13.175 25.574 35.508 30.942 36.806 32.134 6.629 34.717 32.860 29.474 28.923
$F_{DP}$ 18.815 19.006 16.522 18.393 17.007 18.947 17.086 12.524 18.454 16.798 15.004 19.634
$F_{OAE}$ 19.144 2.128 14.899 18.668 16.331 19.294 16.714 6.598 18.249 17.426 15.605 15.099
$F_{EO}$ 0.524 0.193 0.373 0.525 0.442 0.553 0.490 0.147 0.507 0.482 0.447 0.420
Intersection $F_{MEO}$ 16.037 18.196 19.081 16.394 19.895 21.535 13.589 12.135 15.921 14.772 12.424 10.340
$F_{DP}$ 16.749 17.705 16.207 17.525 18.313 19.312 17.029 12.639 17.144 16.813 15.850 16.846
$F_{OAE}$ 7.914 4.565 12.301 6.936 8.704 9.372 5.877 11.469 7.025 5.523 6.909 5.109
$F_{EO}$ 0.381 0.346 0.400 0.399 0.394 0.467 0.317 0.249 0.364 0.336 0.313 0.243
Utility(%) - AUC 0.958 0.967 0.975 0.966 0.964 0.951 0.976 0.978 0.969 0.962 0.967 0.983
ACC 0.864 0.864 0.823 0.876 0.855 0.862 0.909 0.829 0.888 0.874 0.882 0.925
AP 0.934 0.972 0.976 0.948 0.946 0.938 0.964 0.979 0.952 0.942 0.951 0.977
EER 0.087 0.098 0.091 0.082 0.081 0.116 0.064 0.085 0.073 0.084 0.079 0.056
FPR 0.267 0.012 0.349 0.242 0.285 0.272 0.172 0.339 0.221 0.246 0.227 0.142
Table D.7: Detailed fairness and utility evaluation results on a training subset with the ratio of real vs fake is 1:10.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
Xception
 [82]
EfficientB4
 [83]
ViT-B/16
 [84]
F3Net
 [85]
SPSL
 [86]
SRM
 [87]
UCF
 [26]
UnivFD
 [88]
CORE
 [89]
DAW-FDD
 [29]
DAG-FDD
 [29]
PG-FDD
 [30]
Fairness(%) Skin Tone $F_{MEO}$ 6.557 11.357 7.743 7.364 5.850 18.290 5.814 12.493 6.397 5.083 6.515 5.122
$F_{DP}$ 13.087 13.112 12.348 15.751 11.728 22.766 14.881 13.771 14.364 12.265 15.536 12.718
$F_{OAE}$ 1.779 2.939 1.589 1.352 2.402 3.823 1.495 2.386 0.996 1.401 1.471 1.742
$F_{EO}$ 0.108 0.140 0.093 0.153 0.090 0.331 0.135 0.155 0.125 0.084 0.143 0.096
Gender $F_{MEO}$ 1.808 1.803 2.106 1.659 2.316 3.035 3.035 1.997 3.499 1.920 1.812 3.016
$F_{DP}$ 1.769 0.136 0.369 1.321 0.470 2.871 2.503 2.008 2.119 1.097 1.405 2.633
$F_{OAE}$ 1.385 1.269 1.615 1.661 1.777 1.270 1.817 0.579 2.662 1.928 1.734 1.677
$F_{EO}$ 0.025 0.018 0.027 0.032 0.032 0.038 0.035 0.020 0.052 0.036 0.033 0.033
Age $F_{MEO}$ 8.571 11.954 9.832 8.680 9.523 36.509 11.812 10.258 13.180 8.539 9.553 11.109
$F_{DP}$ 17.740 18.338 18.662 17.788 18.910 27.851 16.998 16.446 16.719 18.069 18.034 16.696
$F_{OAE}$ 3.191 0.656 1.405 4.291 2.010 12.139 5.174 2.737 5.723 3.924 4.283 4.859
$F_{EO}$ 0.196 0.157 0.141 0.228 0.174 0.685 0.281 0.173 0.288 0.209 0.240 0.246
Intersection $F_{MEO}$ 11.785 18.146 14.840 14.116 12.940 23.975 11.313 16.240 12.729 10.948 13.399 9.441
$F_{DP}$ 20.192 17.674 19.309 23.135 18.594 29.625 22.394 19.583 21.935 18.674 23.584 17.906
$F_{OAE}$ 3.419 4.233 3.638 4.263 4.020 5.166 3.479 3.133 4.945 3.388 3.885 3.667
$F_{EO}$ 0.251 0.312 0.256 0.329 0.232 0.678 0.305 0.307 0.285 0.225 0.314 0.232
Utility(%) - AUC 0.978 0.973 0.982 0.979 0.982 0.933 0.978 0.975 0.979 0.980 0.982 0.983
ACC 0.920 0.862 0.895 0.928 0.916 0.832 0.921 0.849 0.921 0.920 0.930 0.933
AP 0.978 0.977 0.984 0.978 0.984 0.915 0.979 0.978 0.978 0.979 0.981 0.984
EER 0.070 0.088 0.075 0.066 0.064 0.141 0.076 0.087 0.074 0.070 0.065 0.066
FPR 0.034 0.008 0.009 0.042 0.018 0.116 0.054 0.004 0.055 0.037 0.040 0.051
Table D.8: Detailed fairness and utility evaluation results on a training subset with the ratio of real vs fake is 10:1.
D.2 Performance on Different Age Subgroups
We conduct an analysis of all detectors on age subgroups. 1) As shown in Fig. D.1, facial images with an age range of 0-14 (Child) are more often misclassified as fake, likely due to the underrepresentation of children in our dataset (see Fig. 3 (b)). This suggests detectors tend to show higher error rates for minority groups and show higher accuracy for the majority (Adult). 2) Among those detectors, EfficientB4, UnivFD, and PG-FDD demonstrate a smaller FPR gap between age subgroups, indicating these models may be less susceptible to age bias.
Figure D.1: FPR(%) of each age subgroup. The subgroup with the highest FPR score is highlighted in red, while the subgroup with the lowest FPR score is shown in green.
D.3 Details of Post-Processing
In Section 4 we have applied 6 post-processing methods to evaluate detectors’ robustness. Fig. D.2 visualizes the image after being applied different post-processing methods. We describe each post-processing method as follows:
JPEG Compression: Image compression introduces compression artifacts and reduces the image quality, simulating real-world scenarios where images may be of lower quality or have compression artifacts. In Fig. 6 we apply image compression with quality 80 to each image in the test set.
Gaussian Blur: This post-processing reduces image detail and noise by smoothing it through averaging pixel values with a Gaussian kernel. In Fig. 6 we apply gaussian blur with kernel size 7 to each image in the test set.
Hue Saturation Value: Alters the hue, saturation, and value of the image within specified limits. This post-processing technique is used to simulate variations in color and lighting conditions. Adjusting the hue changes the overall color tone, saturation controls the intensity of colors, and value adjusts the brightness. The results in Fig. 6 are after we adjust hue, saturation, and value with shifting limits 30.
Random Brightness and Contrast: This post-processing method adjusts the brightness and contrast of the image within specified limits. By applying random brightness and contrast variations, it introduces changes in the illumination and contrast levels of the images. This evaluates detector’s robustness to different illumination conditions. The results in Fig. 6 are after we adjust brightness and contrast with shifting limits 0.4.
Random Crop: Resizes the image to a specified size and then randomly crops a portion of it to the target dimensions. This post-processing method is used to evaluate the detector’s robustness to variations in the spatial content of the image. The results in Fig. 6 are after we randomly crop the image with target dimension of $244\times 244$.
Rotation: Rotates the image within a specified angle limit. This post-processing method is used to evaluate the detector’s robustness to changes in the orientation of objects within the image. The results in Fig. 6 are after we randomly rotate the image within a range of -30 to 30 degrees.
Figure D.2: Visualization of the image after different post-processing.
Figure D.3: Robustness analysis in terms of utility and fairness under varying degrees of JPEG compression.
Figure D.4: Robustness analysis in terms of utility and fairness under varying kernel sizes of Gaussian Blur.
Figure D.5: Robustness analysis in terms of utility and fairness under varying degrees of Hue Saturation Value.
Figure D.6: Robustness analysis in terms of utility and fairness under varying degrees of Rotations.
Figure D.7: Robustness analysis in terms of utility and fairness under varying degrees of Brightness Contrast.
D.4 Additional Fairness Robustness Evaluation Results
Fig. D.3 to Fig. D.7 demonstrate detectors’ robustness analysis in more detail as a function of different degrees of post-processing. Overall, ViT-B/16 [84] and UnivFD [88] show stronger robustness to various post-processing methods compared to other detection methods. Fairness-enhanced detectors do not have robustness against post-processing; this would be a direction for future studies to work on. Figure D.3 presents a detailed robustness analysis in terms of utility and fairness under varying degrees of JPEG compression. The utility of all detectors decreases as image quality is reduced. Among the detectors, ViT-B/16 [84] exhibits the highest utility robustness, ViT-B/16 [84] and UnivFD [88] both demonstrate the strongest fairness robustness. When considering Gaussian blur, ViT-B/16 again stands out as the most robust detector in terms of utility, whereas DAW-FDD [29] and UnivFD [88] show the great robustness in terms of fairness. Against Hue Saturation Value adjustments, SPSL [86] shows the strongest utility robustness, while the fairness of DAW-FDD [29] fluctuates less with different Hue Saturation Value adjustments. ViT-B/16 demonstrates superior robustness in both utility and fairness when facing rotations. For brightness contrast variations, SPSL [86] is the most robust detector in terms of utility, while UnivFD once again shows superior robustness in terms of fairness. Last, we can get the same conclusion from Fig. D.3 to Fig. D.7 as in the main manuscript, that post-processing clearly impairs detectors’ utility but does not necessarily make detectors more biased.
D.5 Full Results of Effect of Increasing the Size of Train Set
In this section, we provide the full evaluation results tested under different sizes of train set, as shown from Table D.2 to Table D.5. Intersection $F_{EO}$ and AUC align with the results in Fig. 7 of the submitted manuscript.
D.6 Full Results of Effect of the Ratio of Real and Fake
In this section, we provide the full evaluation results tested under the train set with different ratios of real and fake, as shown from Table D.6 to Table D.8. Intersection $F_{EO}$ and AUC align with the results in Fig. 7 of the submitted manuscript.
D.7 Comparison Results with Foundation Model
In the Discussion (Section 5.3) of the main manuscript, we highlighted the potential of integrating foundation models (e.g., CLIP) into detector design as a strategy for mitigating bias. To explore this, we conducted a preliminary experiment by designing a detector using a frozen CLIP model combined with a trainable 3-layer MLP. This model was trained and tested on the AI-Face dataset. For comparison, we selected one representative detector from each model type: EfficientB4 [83], SPSL [86], UnivFD [88], and PG-FDD [30]. These four detectors’ results are consistent with those reported in Table 4. As shown in Table D.9, the CLIP+MLP detector demonstrates a clear advantage in both fairness and utility metrics, suggesting that foundation models hold significant promise for bias mitigation. For instance, its $F_{EO}$ score is 3.11% lower than the second-best method, PG-FDD, for the Skin Tone group, and 14.046% lower for the Intersection group.
Fairness Utiliy
Skin Tone Gender Age Intersection -
Method $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ AUC ACC AP EER FPR
EfficientB4 5.385 1.725 1.487 5.863 8.300 6.184 4.377 11.062 6.796 11.849 2.856 10.300 17.586 8.607 8.461 25.114 98.611 94.203 99.542 6.689 20.066
SPSL 4.411 1.827 1.037 4.534 8.055 9.379 1.135 9.789 27.614 11.232 7.270 40.943 10.379 13.259 2.464 21.679 98.747 96.346 99.356 4.371 13.661
UnivFD 4.503 1.19 1.622 5.408 2.577 8.556 2.748 5.536 5.436 15.249 3.793 14.148 6.119 14.026 6.287 20.255 98.192 93.651 99.400 7.633 18.550
PG-FDD 3.190 1.252 1.071 3.702 6.465 9.746 0.882 9.115 14.804 10.467 5.009 29.585 9.578 14.697 3.062 18.348 99.172 96.174 99.694 4.961 10.971
CLIP+MLP 0.419 0.938 0.227 0.591 0.506 8.658 0.334 1.021 0.765 14.473 0.395 1.802 1.973 13.992 1.000 4.302 99.973 99.290 99.991 0.793 1.171
Table D.9: Fairness and utility performance of CLIP+MLP compared to representative detectors on the AI-Face dataset, highlighting the potential of foundation models for bias mitigation.
Appendix E Datasheet for AI-Face
In this section, we present a DataSheet [114] for AI-Face.
E.1 Motivation For Dataset Creation
  • Why is the dataset created? For researchers to evaluate the fairness of AI face detection models or to train fairer models. Please see Section 2 ‘Background and Motivation’ in the submitted manuscript.
  • Has the dataset been used already? Yes. Our fairness benchmark is based on this dataset.
  • What (other) tasks could the dataset be used for? Could be used as training data for generative methods attribution task.
E.2 Data Composition
  • What are the instances? The instances that we consider in this work are real face images and AI-generated face images from public datasets.
  • How many instances are there? We include 1,646,545 face images from public datasets. Please see Table B.1 for details.
  • What data does each instance consist of? Each instance consists of an image.
  • Is there a label or target associated with each instance? Each image is associated with gender annotation, age annotation, skin tone annotation, intersectional attribute (gender and skin tone) annotation, and target label (fake or real).
  • Is any information missing from individual instances? No.
  • Are relationships between individual instances made explicit? Not applicable – we do not study the relationship between each image.
  • Does the dataset contain all possible instances or is it a sample? Contains all instances our curation pipeline collected. Since the current dataset does not cover all available images online, there is a high probability more instances can be collected in the future.
  • Are there recommended data splits (e.g., training, development/validation, testing)? For detector development and training, the dataset can be split as 6:2:2.
  • Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description. Yes. Despite our extensive efforts to mitigate the bias that may introduced by the automated annotator and reduce demographic label noise, there may still be mislabeled instances. Given the dataset’s size of over 1 million images and most are generated face images, it is impractical for humans to manually check and correct each image individually.
  • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? The dataset is self-contained.
E.3 Collection Process
  • What mechanisms or procedures were used to collect the data? We build our AI-Face dataset by collecting and integrating public AI-generated face images sourced from academic publications, GitHub repositories, and commercial tools. Please see ‘Data Collection’ in Section 3.1
  • How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data? The data can be acquired after our verification of user submitted and signed EULA.
  • If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)? Not applicable. We did not sample data from a larger set. But we use RetinaFace [66] for detecting and cropping faces to ensure each image only contains one face.
  • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. The data was collected from February 2024 to April 2024, even though the data were originally released before this time. Please refer to the cited papers in Table B.1 for specific original data released time.
E.4 Data Processing
  • Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? Yes. We discussed in ‘Data Collection’ in Section 3.1.
  • Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the ‘raw’ data. The ‘raw’ data can be acquired through the original data publisher. Please see the cited papers in Table B.1.
  • Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point. Yes. We use RetinaFace [66] for detecting and cropping faces to ensure each image only contains one face. Demographic annotations are given by our annotator, see ‘Annotation Generation’ in Section 3.2. Our annotator code will not be released considering the ethical guidelines.
  • Does this dataset collection/processing procedure achieve the motivation for creating the dataset stated in the first section of this datasheet? If not, what are the limitations? Yes. The dataset does allow for the study of our goal, as it covers comprehensive generation methods, demographic annotations for evaluating current detectors and training fairer detectors.
E.5 Dataset Distribution
  • How will the dataset be distributed? We distribute all the data as well as CSV files that formatted all annotations of images under the CC BY-NC-ND 4.0 license and strictly for research purposes.
  • When will the dataset be released/first distributed? What license (if any) is it distributed under? The dataset will be released following the paper’s acceptance, and it will be under the permissible CC BY-NC-ND 4.0 license for research-based use only. Users can access our dataset by submitting an EULA.
  • Are there any copyrights on the data? We believe our use is ‘fair use’ since all data in our dataset is collected from public datasets.
  • Are there any fees or access restrictions? No.
E.6 Dataset Maintenance
  • Who is supporting/hosting/maintaining the dataset? The first author of this paper.
  • Will the dataset be updated? If so, how often and by whom? We do not plan to update it at this time.
  • Is there a repository to link to any/all papers/systems that use this dataset? Our fairness benchmark uses this dataset, a brief instruction of how to use this dataset and the code of fairness benchmark is on https://github.com/Purdue-M2/AI-Face-FairnessBench.
  • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? Not at this time.
E.7 Legal and Ethical Considerations
  • Were any ethical review processes conducted (e.g., by an institutional review board)? No official processes were done since all data in our dataset were collected from the existing public datasets.
  • Does the dataset contain data that might be considered confidential? No. We only use data from public datasets.
  • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why No. It is a face image dataset, we have not seen any instance of offensive or abusive content.
  • Does the dataset relate to people? Yes. It is a face image dataset containing real face images and AI-generated face images.
  • Does the dataset identify any subpopulations (e.g., by age, gender)? Yes, through demographic annotations.
  • Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? Yes. It is a face image dataset. The age, gender, and skin tone can be identified through the face image, also through the demographic annotation we provide. All of the images that we use are from publicly available data.
E.8 Author Statement and Confirmation of Data License
The authors of this work declare that the dataset described and provided has been collected, processed, and made available with full adherence to all applicable ethical guidelines and regulations. We accept full responsibility for any violations of rights or ethical guidelines that may arise from the use of this dataset. We also confirm that the dataset is released under the CC BY-NC-ND 4.0 license, permitting sharing and downloading of the work in any medium, provided the original author is credited, and it is used non-commercially with no derivative works created.